Yohkoh Analysis Guide

Reference Guide

Ver 1.06 Ed: M.D. Morrison 6-Nov-92 Ver 2.00 Ed: M.D. Morrison 25-Jan-94 Ver 2.91 Ed: R.D. Bentley 1-Mar-98

This Copy Produced: March 6, 1998

Prepared at Mullard Space Science Laboratory University College London

Contents

1 Introduction to the Yohkoh Data and Software 1 1.1 YS, YDB, and YDn Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Yohkoh Software Organization . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 How to Contribute Software to the Yohkoh Tree . . . . . . . . . . . . . . . . . . 2 1.4 Yohkoh Data File Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.5 Yohkoh Data File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1.5.1 Yohkoh Directory Logicals/Environment Variables . . . . . . . . . . . . . 4 1.5.2 Yohkoh Prefixes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.5.3 Yohkoh File IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.5.4 Yohkoh Week IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.5.5 Yohkoh Carrington IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

1.6 Interactive Data Language (IDL) . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.6.1 Some Comments about IDL . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.6.2 Programs, Procedures, and Functions . . . . . . . . . . . . . . . . . . . . 10 1.6.3 Running IDL programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

2 General Yohkoh Software 12

2.1 Routines for Reading and Saving Data . . . . . . . . . . . . . . . . . . . . . . . . 12 2.1.1 RD XDA, RD xxx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.1.2 RD ROADMAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.1.3 RD QS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.1.4 SAVEGEN and RESTGEN . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.1.5 SAV SDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.1.6 SAV BDA and SAV BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.1.7 SAV HXI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.1.8 SXT2FITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.2 Routines for Data Selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2.2.1 TIM2DSET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.3 Routines for Making Time Plots . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2.3.1 PLOTY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.3.2 OBS PlOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.4 Routines for Data and Information Extraction . . . . . . . . . . . . . . . . . . . . 15 2.4.1 The Concept Behind the GT Routines . . . . . . . . . . . . . . . . . . . . 15 2.4.2 GT DP RATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.4.3 DPRATE2SEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.4.4 GT DP MODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.4.5 GET INFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

2.5 Routines for Calibration and Analysis . . . . . . . . . . . . . . . . . . . . . . . . 18 2.5.1 FSP PROC [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.5.2 FSP 11 [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

2.6 Routines for Image Alignment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ii CONTENTS

2.6.1 SXT PREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.6.2 Other Routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

2.7 Accessing Secondary Databases . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.7.1 CONTACTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.7.2 PR FEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.7.3 PR EVN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.7.4 PLOT EVN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.7.5 RD FEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.7.6 RD OBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

2.8 Reading Exabyte Archive Tapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3 Bragg Crystal Spectrometer (BCS) 22

3.1 Routines for Data Selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.1.1 LIST BDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.1.2 SELECT BDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

3.2 Routines for Making Time Plots . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.2.1 BCS 24HR PLOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.2.2 PLOTT BDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.2.3 LCBDA [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

3.3 Routines for Making Spectral Plots . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.3.1 PLOTS BDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.3.2 PLOTBDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.3.3 BCS MULTI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.3.4 BCS CONT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

3.4 Routines for 2-D Spectral Display . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.4.1 DISP BDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.4.2 BCS SPMOVIE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.4.3 GS [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

3.5 Routines for Data and Information Extraction . . . . . . . . . . . . . . . . . . . . 24 3.5.1 EXT BCSCHAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.5.2 LIST BDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.5.3 GT TOTAL CNTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.5.4 GT BLOCKID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

3.6 Routines for Calibration and Analysis . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.1 BCS DECOMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.2 BCS NORM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.3 SUM BDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.4 PLOT REF [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.5 BCS DTFAC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.6 BCS POINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.7 BCS SYNTHETIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.6.8 BCS SPEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.7 BSC Routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

3.7.1 HELP BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.7.2 MK BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.7.3 SEL BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.7.4 LIST BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.7.5 SAV BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.7.6 RD BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.7.7 BSC2BSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.7.8 MAP BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.7.9 LC BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.7.10 WBDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.7.11 GT BSC BINCAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.7.12 FIT BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.7.13 BCS BROAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 3.7.14 PLOT BSC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 3.7.15 FIT BSC PLOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 3.7.16 MK FIT BSC EPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

4 Hard X-Ray Telescope (HXT) 36

4.1 Routines for Making Time Plots . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

4.1.1 PLOTT HDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.2 Routines for 2-D Spectral Display . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

4.2.1 DISP HDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.3 Routines for Data and Information Extraction . . . . . . . . . . . . . . . . . . . . 36

4.3.1 GT SUM L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.3.2 GET INFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

4.4 Routines for Calibration and Analysis . . . . . . . . . . . . . . . . . . . . . . . . 36 4.4.1 HXT QLOOK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.4.2 HXTPRO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.4.3 HXT IMG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.4.4 GET HXT POS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.4.5 HXT FSP and HXTBOX FSP [*] . . . . . . . . . . . . . . . . . . . . . . 38 4.4.6 SXTHXT FSP and SXTHXTBOX FSP[*] . . . . . . . . . . . . . . . . . . 38 4.4.7 HXT DECOMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

5 Soft X-Ray Telescope (SXT) 40

5.1 Routines for Data Selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.1.1 SSWHERE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.1.2 SHOW OBS3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.1.3 SHOW OBS4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.1.4 PLOT FOV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

5.2 Routines for Making Time Plots . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 5.2.1 LCUR IMAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 iv CONTENTS

5.3 Routines for Image Display and Enhancement . . . . . . . . . . . . . . . . . . . . 42 5.3.1 STEPPER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 5.3.2 XY RASTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 5.3.3 DISP MONTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 5.3.4 UNSHARP MASK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 5.3.5 DE SPIKER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.3.6 SXT GRID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

5.4 Routines for Data and Information Extraction . . . . . . . . . . . . . . . . . . . . 44 5.4.1 GET INFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.4.2 EXT SUBSET [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.4.3 GT FILTA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.4.4 GT FILTB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.4.5 GT PFI FFI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.4.6 GT RES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.4.7 GT COMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.4.8 GT DPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.4.9 GT MBE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.4.10 GT EXPDUR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.4.11 GT EXPLAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 5.4.12 GT EXPMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 5.4.13 GT CORNER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 5.4.14 GT CENTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 5.4.15 GT SXT AXIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 5.4.16 GT PIX SIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 5.4.17 GT TEMP CCD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.4.18 GT TEMP HK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.4.19 SXTPNT SUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.4.20 PR SXTOBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.4.21 PFI LOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

5.5 Routines for Calibration and Analysis . . . . . . . . . . . . . . . . . . . . . . . . 51 5.5.1 SXT PREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 5.5.2 XSXT PREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 5.5.3 SXT DECOMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 5.5.4 SXT COMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 5.5.5 RESTORE LOW8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 5.5.6 SFD COMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 5.5.7 SFD DECOMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 5.5.8 EXP NORM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 5.5.9 SXT COMPOSITE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5.5.10 MK SFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5.5.11 DARK SUB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.5.12 GET DC IMAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.5.13 LEAK SUB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.5.14 INTERP IMG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.5.15 SXT FSP and SXTBOX FSP [*] . . . . . . . . . . . . . . . . . . . . . . . 57 5.5.16 SXTHXT FSP and SXTHXTBOX FSP [*] . . . . . . . . . . . . . . . . . 58 5.5.17 SXT TEEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.5.18 SXT ERG PER DN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.5.19 SXT VIGNETTE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.5.20 SXT OFF AXIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.5.21 WL PROCESS [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

5.6 Routines for Image Alignment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.6.1 SXT PREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.6.2 SXT CENTER [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.6.3 GET GBO PFI [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.6.4 OCENTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

5.7 Routines for Accessing Secondary Database Files . . . . . . . . . . . . . . . . . . 61 5.7.1 GTAB COMM, GTAB ENTRY, GTAB ROI . . . . . . . . . . . . . . . . 61 5.7.2 GTAB PFI and GTAB FFI . . . . . . . . . . . . . . . . . . . . . . . . . . 62

5.8 Routines for Engineering Plots . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 5.8.1 PLOT SOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 5.8.2 PLOT TEMPS2 [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 5.8.3 PLOT SSL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

6 Wide Band Spectrometer (WBS) 64

6.1 Routines for Making Time Plots . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

6.1.1 PLOTT WDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 6.2 Routines for Making Spectral Plots . . . . . . . . . . . . . . . . . . . . . . . . . . 64

6.2.1 PLOTS WDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 6.3 Routines for 2-D Spectral Display . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

6.3.1 DISP WDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 6.3.2 SXSPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 6.3.3 WBSPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.3.4 PLOT HXSPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.3.5 PLOT HXSPH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.3.6 PLOT GRSPCL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.3.7 PLOT GRSPCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.3.8 PLOT GRSPHL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.3.9 PLOT GRSPHH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 6.3.10 Light Curve of RBM-PC . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

6.4 Routines for Data and Information Extraction . . . . . . . . . . . . . . . . . . . . 66 6.4.1 GT Routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 6.4.2 GET INFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

6.5 Routines for Calibration and Analysis . . . . . . . . . . . . . . . . . . . . . . . . 66 6.5.1 HXS FSP [*] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 6.5.2 HXS SPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 6.5.3 HXS SP2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 6.5.4 GRS SPEFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 6.5.5 Further Analysis of HXS and GRS Data . . . . . . . . . . . . . . . . . . . 68

6.6 Instrument Response Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 6.6.1 SXS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 6.6.2 HXS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 6.6.3 GRS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

A Yohkoh Distribution and Archive Tapes 71 A.1 The Organization of the Tapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 A.2 Overview of Yohkoh Data Archive . . . . . . . . . . . . . . . . . . . . . . . . . . 71 A.3 Structure of Archive Tape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 A.4 Access Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

A.4.1 Reading Tapes (RDTAP) . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 A.4.2 Tape Reading on VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 A.4.3 Tape Read Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 A.4.4 Writing Tapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

B Institutes and Nodes with the Yohkoh Package 77 B.1 List of Institutes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 B.2 List of Node Names and Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . 78

C Accessing the University of Hawaii SPAM Database 81

D Accessing the Optical Solar Data of NAOJ 82

E Using the Laser Disk System at ISAS 90 E.1 Viewing Existing Movies (VIDEO MENU) . . . . . . . . . . . . . . . . . . . . . . 90 E.2 Making Title Pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 E.3 Recording a New Movie to Laser Disk . . . . . . . . . . . . . . . . . . . . . . . . 92

E.3.1 Hardware: Positioning the Sony Laser Disk . . . . . . . . . . . . . . . . . 92 E.3.2 Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 E.3.3 Adding Another Movie . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 E.3.4 NVS/Sony Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

E.4 Transferring a Movie to VCR Video . . . . . . . . . . . . . . . . . . . . . . . . . 93 E.4.1 Initial setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 E.4.2 Short test of hardware communications. . . . . . . . . . . . . . . . . . . . 94 E.4.3 Running MK VCR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 E.4.4 How to create/modify the file that MK VCR uses . . . . . . . . . . . . . 95 F Using the Laser Disk System at LPARL 97

F.1 Viewing Existing Movies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 F.1.1 Using IDL and Menu Driven Software . . . . . . . . . . . . . . . . . . . . 97 F.1.2 Displaying a Single Recorded Image . . . . . . . . . . . . . . . . . . . . . 97 F.1.3 Playing the Panasonic Laser Disks . . . . . . . . . . . . . . . . . . . . . . 97

F.2 Making Title Pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 F.3 Recording a New Movie to Laser Disk . . . . . . . . . . . . . . . . . . . . . . . . 98

F.3.1 Panasonic Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 F.4 Transferring a Movie to VCR Video . . . . . . . . . . . . . . . . . . . . . . . . . 100

G Using Magneto-Optical (MO) Disks at ISAS 101 G.1 Mounting MO Disks From Unix . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 G.2 Mounting MO Disks From IDL . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

H Guides to Yohkoh Operations at ISAS 102 viii LIST OF FIGURES

List of Figures 1 Introduction to the Yohkoh Data and Software

1.1 YS, YDB, and YDn Directories

The Yohkoh data are organized into three general parts.

YS The software, documentation, and some calibration data files are located in the Yohkoh Software (YS) tree. There is a separate branch for each instrument, and one for general software which holds software and data that can be used by all instruments.

YDB The secondary database files are contained in the Yohkoh Database (YDB) tree. The secondary database files include S/C attitude information, observing logs, event logs, detailed calibration files, summary data files, and several ground-based database files.

YD The raw reformatted Yohkoh data (YD) are placed in the YD directories. At ISAS, there are twenty one Yohkoh data directories (/yd0, /yd1, /yd2, .... /yd20). The standard method for putting data on-line is to create a directory with the week number, for example /yd10/93 41a and put all the data for week 93 41 into that directory (except for the weekly files).

1.2 Yohkoh Software Organization

It is assumed that you have successfully installed the Yohkoh software on your system or you are using a machine that has the Yohkoh software. In addition to having made the software installation, it is necessary to have executed the Yohkoh initialization routine (‘/ys/gen/script/idl setup’ or ‘.yslogin’ for the Unix machines). If you have not done this yet, see the Appendix of the Yohkoh Reference Guide.

The Yohkoh software is organized under one tree. The top directory is $ys for Unix systems, and is the logical YS: for the VMS systems. Under that tree are the following branches:

Branch  Description
SITE  Site specific software
GEN  General software and documents that all instruments can use
BCS F77 BCS FORTRAN specific software and documentation
BCS  BCS specific software and documentation
HXT  HXT specific software and documentation
HXT F77 HXT FORTRAN specific software and documentation
SXT  SXT specific software and documentation
WBS  WBS specific software and documentation
ATEST Newly created or modified software and documentation
UCON  User Contributed software and documentation
IDLFIX Work-arounds for systems using older IDL versions

Under each of the above instrument directories, there are the following directories (for example, ‘/ys/sxt/doc’)

Branch  Description
DOC  Documentation
RESPONSE Instrument calibration and response data files
SOFT  Software
STATUS  Instrument status information

Software that has been thoroughly tested is put into the instrument release directory ‘soft’.^{1} Software which has just been written, or software that is modified is put into the ‘atest’ area for a period of a few weeks. Certain privileged users have write privilege to those directories. If problems develop with modified software, it is possible to recover the old version by copying it from the ‘soft’ directory. The software developed by general users is placed online under the ‘ucon’ (user contributed) branch. Generally each user who is contributing has a directory of their own. The directories under the ‘soft’ branch are broken up by function for the ‘gen’ and instrument branchs, but by person under the ‘ucon’ branch. If a user cannot remember the name of a function, he/she can do a listing on the different directories and will probably recognize it. A list of the directories that currently exist for the instrument teams is:

GEN  BCS  HXT SXT  WBS
dbase  bda  util register  util
gbo  bsc  sensitivity
jhuapl  bsd  util
mo disk to be deleted  widgets
movie  util

orbit pointing ref access reformat spectra tape util utplot

1.3 How to Contribute Software to the Yohkoh Tree

See the appendix of the Reference Guide.

1The previous organization had an ‘atest’, ‘rel’, and ‘usercontrib’ branch under each ‘soft’ directory.

SOLAR-A reformatted files Road Map Section Optional Data Section Data Data Index  Index and Data Section Data Data Index Data Data Index  Quasi-Static Index Section File Header File Info/Pointer Section

1.4 Yohkoh Data File Format

For a full description of the Yohkoh database and all the different data structures, see the File Control Document.

Each file is logically divided into the following six sections. Some files will not use all the sections described below, but all will have a Pointer and File Header Section.

1. File Information and Pointer Section

2. File Header Section

3. Quasi-Static Index Section

4. Index and Data Section

5. Optional Data Section

6. Road Map Section

The program which reads the file learns from the Pointer Section how to read the rest of the file and where to go to get certain data.

The File Header Section provides information on what data are contained in the file, generally the extent of the time covered by the contents.

The Quasi-Static Section of the file contains index information that does not vary during the course of an orbit, or varies slowly.

The Index and Data section contains ‘data sets’. A data set is a single image for SXT, single spectrum for BCS, a single major frame of data for HXT, and two major frames of data for WBS (it takes two major frames for a complete set of WBS data). For each data set there is an index which describes the date and time that the data were taken, the mode and position of the instrument’s peripherals (e.g., filters, HV), and information on temperature and gain information.

The Optional Data Section is only used by the BCS data (BDA) files and the spacecraft attitude (ADA) files. The BDA files hold the ‘DP SYNC’ information, which is information that is coming down every major frame. Since the BCS spectra are asynchronous to the major frames, it is stored separately. The ADA file holds the full 2048 point HXA scans.

The Roadmap Section allows a user to access a brief summary of the contents of the file and to perform searches on that summary to select what data should be extracted.

1.5 Yohkoh Data File Names

1.5.1 Yohkoh Directory Logicals/Environment Variables

There is a set of VMS logicals and Unix environment variables which point to the database directories. By using these logicals in all of the access routines (along with the routine CONCAT DIR) it is possible for software to be directly portable between VMS and Unix machines. The following is a list of the logicals that exist for the Yohkoh database.

DIR GEN ADS  DIR SXT ATABLES  DIR BCS ATODAT
DIR GEN ATR  DIR SXT CAL  DIR BCS BALDAT
DIR GEN ATT  DIR SXT CALIBRATE DIR BCS CAL
DIR GEN DATA  DIR SXT DOC  DIR BCS CALDAT
DIR GEN DOC  DIR SXT ENGIN  DIR BCS CAT
DIR GEN EVN  DIR SXT SDC  DIR BCS DOC
DIR GEN FEM  DIR SXT SDL  DIR BCS EXE
DIR GEN GBE  DIR SXT SDW  DIR BCS LOGS
DIR GEN GBL  DIR SXT SENSITIVE DIR BCS MICRO
DIR GEN GEV DIR SXT SFC DIR BCS SYNSPEC
DIR GEN GOL  DIR SXT SFM
DIR GEN GXT DIR SXT SFS DIR GBO TEMP
DIR GEN MO  DIR SXT SLD
DIR GEN MOVIE  DIR SXT SLS
DIR GEN NAR  DIR SXT SOT
DIR GEN NEL  DIR SXT SSL
DIR GEN NTS  DIR SXT SSX
DIR GEN OBD  DIR SXT SXA
DIR GEN OBS  DIR SXT SXC
DIR GEN ORBIT  DIR SXT SXL
DIR GEN ORBIT RAW  DIR SXT SXL2
DIR GEN ORBIT SOL  DIR SXT TABLES

DIR GEN ORBIT SW

DIR GEN OSF  DIR WBS CAL

DIR GEN OSP

DIR GEN OWH  DIR HXT CAL
DIR GEN PAN LASER  DIR HXT EXE

DIR GEN PNT DIR GEN SCRIPT DIR GEN SETUP DIR GEN SETUPD DIR GEN SETUPDBACK DIR GEN SPECTRA DIR GEN STATUS DIR GEN SYNOPTIC DIR GEN TAPECOPY DIR GEN TBB DIR GEN XAD DIR GEN XBD 6 1. INTRODUCTION TO THE YOHKOH DATA AND SOFTWARE 1.5.2 Yohkoh Prefixes

Prefix Description  File Type
ADA S/C Attitude Raw Reformatted Data  per Orbit
ATR Raw S/C Attitude Database  Weekly
ATT Processed S/C Attitude Database  Weekly
BDA BCS Raw Reformatted Data  per Orbit
BSD *OLD* BCS Instrument Calibrated Spectra  User Specified
BPC *OLD* BSDCAL Output (parameters)  User Specified
BTH *OLD* BSDFIT Output (Fitted Spectra)  User Specified
BFT *OLD* BSDFIT Output (parameters)  User Specified
BSC *NEW* BCS Instrument Calibrated Spectra  User Specified
BSA *NEW* Answer File for BSC Generation  User Specified
BSF *NEW* Output from Spectral Fitting  User Specified
CBA S/C Common Basic Raw Reformatted Data  per Orbit
EVN Yohkoh Event Log  Weekly
FEM Yohkoh Orbital Ephemeris  Weekly
GEV GOES Event Log  Weekly
GXT GOES One Minute Light Curve Data  Weekly
G6D GOES 3 Sec Light Curve for S/C 6  Weekly
G7D GOES 3 Sec Light Curve for S/C 7  Weekly
GXD GOES Derived One Minute Light Curve Data  Weekly
GBE GRO BATSE Event Log  Mission
GBL GRO BATSE Light Curves  Weekly
G x Ground-Based Observation (GBO) Images (Sites) per Image (FITS)
Gx  Ground-Based Observation (GBO) Images (Data Type) per Image (FITS)
GBC GBO Big Bear Large Scale H-alpha PFI FITS Image  per Image (FITS)
GBH GBO Big Bear H-alpha FITS Image  per Image (FITS)
GBK GBO Big Bear Calcium FITS Image  per Image (FITS)
GBW GBO Big Bear White Light FITS Image  per Image (FITS)
GCH GBO Boulder Colorado H-alpha FITS Image  per Image (FITS)
GGH GBO GSFC H-alpha FITS Image  per Image (FITS)
GHH GBO Holloman H-alpha FITS image  per Image (FITS)
GLH GBO Leamonth Australia H-Alpha FITS Image  per Image (FITS)
GNH GBO NAOJ Japan H-alpha FITS Image  per Image (FITS)
GJH GBO CRL Hiraiso H-alpha FITS Image  per Image (FITS)
GKI GBO Kitt Peak He 10830 FITS Image  per Image (FITS)
GKM GBO Kitt Peak Magnetogram FITS Image  per Image (FITS)
GNP GBO Nobeyama Polarized Radio Image (17 GHz)  per Image (FITS)
GNT GBO Nobeyama Total Intensity Radio Image (17 GHz) per Image (FITS)
GOL GBO Observing Log  Weekly

Prefix Description  File Type
HDA HXT Raw Reformatted Data  per Orbit
HXI HXT Synthesized Image  User Specified
MDM Magneto-Optical (MO) Map  Mission
NAR NOAA Active Region  Weekly
NEL Nobeyama Radio Observatory Event Log  Mission
NTS Nobeyama Radio Observatory Time Series Weekly
OBS *OLD* Yohkoh Observing Log  Weekly
OBD BCS Observing Log  Weekly
OSF SXT Full Frame Observing Log  Weekly
OSP SXT Partial Frame Observing Log  Weekly
OWH WBS/HXT Observing Log  Weekly
PNT *OLD* S/C Pointing File  Weekly
SDL SXT Dark Current Log  Weekly
SDC SXT Dark Current Images  Weekly (SDA)
SDW SXT Dark Current Images for Warm CCD Weekly (SDA)
SDP SXT Dark Current PFI Images  Weekly (SDA)
SFD SXT FFI Desaturated Composite Images  Weekly (SDA)
SFR SXT FFI Raw Reformatted Data  per Orbit (SDA)
SFS SXT FFI Special Images (diffuser, flood...) Weekly (SDA)
SFT SXT FFI Terminator Images  Weekly (SDA)
SFW SXT FFI White Light Images  Weekly (SDA)
SFC SXT FFI Calibration (Leak) Image  per Image (SDA)
SFM SXT FFI Monthly File (daily SFD image) Monthly (SDA)
SLS SXT Leak Scatter Log  Weekly
SLD SXT Leak Dark Current Log  Weekly
SPR SXT PFI Raw Reformatted Data  per Orbit (SDA)
SOT SXT Optical Telescope Log  Weekly
SSC SXT Synoptic Images (centered)  Weekly (SDA)
SSE SXT Synoptic Images (east of center)  Weekly (SDA)
SSW SXT Synoptic Images (west of center)  Weekly (SDA)
SSL SXT Summary Log  Weekly
SXC SXT X-Ray Sun Center Log  Weekly
SXL SXT X-Ray Histogram Log  Weekly
SXA SXT X-Ray Average  Weekly
WDA WBS Raw Reformatted Data  per Orbit
XAD Exabyte ASCII Directory for Archive Tape Weekly per Tape
XBD Exabyte Binary Directory for Archive Tape Weekly per Tape

1.5.3 Yohkoh File IDs The orbit file ID’s are 11 characters long in the format shown below:

        YYMMDD.HHMM
        where YY - Year of data
              MM - Month of data
              DD - Day of data
              HH - Hours
              MM - Minutes

1.5.4 Yohkoh Week IDs

The Yohkoh week starts on Sunday and ends on Saturday. CAUTION: Yohkoh data taken at 00:01 (one minute after midnight) on Sunday morning is most likely in a file and tape for the previous week since the orbit started on Saturday and the first full orbit for Sunday has not started. Up to the first 90 minutes of data taken on Sunday could be stored in the previous week, depending on when the spacecraft day begins.

The weekly IDs are of the form:

        YY_WWa.NN
        where YY - Year of data
              WW - Week number of the data (1 to 53)
               a - is fixed (reserved for future use)
              NN - is the program version number which created the file

A list of the dates covered by each of the weeks is given in Appendix xxx.

1.5.5 Yohkoh Carrington IDs

The Carrington IDs are of the form:

        _crRRRRa.NN
      where cr   - signifies a Carrington Rotation ID
            RRRR - is the rotation number
             a   - is fixed (reserved for future use)
            NN   - is the program version number which created the file

The dates for the Carrington Rotations are defined in the table.

1.6 Interactive Data Language (IDL)

Most of the Yohkoh software is written in the programming language IDL of Research Systems, Inc (RSI). If certain rules are followed, then the same IDL software can be run on a wide variety of computers.

Carrington Starting Date Carrington Starting Date Carrington Starting Date
Rotation  Rotation  Rotation

1846  21-Aug-91  1888  10-Oct-94  1930  29-Nov-97
1847  18-Sep-91  1889  6-Nov-94  1931  26-Dec-97
1848  15-Oct-91  1890  3-Dec-94  1932  22-Jan-98
1849  11-Nov-91  1891  31-Dec-94  1933  18-Feb-98
1850  8-Dec-91  1892  27-Jan-95  1934  18-Mar-98
1851  5-Jan-92  1893  23-Feb-95  1935  14-Apr-98
1852  1-Feb-92  1894  23-Mar-95  1936  11-May-98
1853  28-Feb-92  1895  19-Apr-95  1937  7-Jun-98
1854  26-Mar-92  1896  16-May-95  1938  5-Jul-98
1855  23-Apr-92  1897  12-Jun-95  1939  1-Aug-98
1856  20-May-92  1898  10-Jul-95  1940  28-Aug-98
1857  16-Jun-92  1899  6-Aug-95  1941  25-Sep-98
1858  14-Jul-92  1900  2-Sep-95  1942  22-Oct-98
1859  10-Aug-92  1901  29-Sep-95  1943  18-Nov-98
1860  6-Sep-92  1902  27-Oct-95  1944  15-Dec-98
1861  3-Oct-92  1903  23-Nov-95  1945  12-Jan-99
1862  31-Oct-92  1904  20-Dec-95  1946  8-Feb-99
1863  27-Nov-92  1905  17-Jan-96  1947  7-Mar-99
1864  24-Dec-92  1906  13-Feb-96  1948  3-Apr-99
1865  20-Jan-93  1907  11-Mar-96  1949  1-May-99
1866  17-Feb-93  1908  7-Apr-96  1950  28-May-99
1867  16-Mar-93  1909  5-May-96  1951  24-Jun-99
1868  12-Apr-93  1910  1-Jun-96  1952  22-Jul-99
1869  10-May-93  1911  28-Jun-96  1953  18-Aug-99
1870  6-Jun-93  1912  26-Jul-96  1954  14-Sep-99
1871  3-Jul-93  1913  22-Aug-96  1955  11-Oct-99
1872  30-Jul-93  1914  18-Sep-96  1956  8-Nov-99
1873  27-Aug-93  1915  15-Oct-96  1957  5-Dec-99
1874  23-Sep-93  1916  12-Nov-96  1958  1-Jan-00
1875  20-Oct-93  1917  9-Dec-96  1959  29-Jan-00
1876  17-Nov-93  1918  5-Jan-97  1960  25-Feb-00
1877  14-Dec-93  1919  1-Feb-97  1961  23-Mar-00
1878  10-Jan-94  1920  1-Mar-97  1962  19-Apr-00
1879  6-Feb-94  1921  28-Mar-97  1963  17-May-00
1880  6-Mar-94  1922  24-Apr-97  1964  13-Jun-00
1881  2-Apr-94  1923  22-May-97  1965  10-Jul-00
1882  29-Apr-94  1924  18-Jun-97  1966  6-Aug-00
1883  26-May-94  1925  15-Jul-97  1967  3-Sep-00
1884  23-Jun-94  1926  11-Aug-97  1968  30-Sep-00
1885  20-Jul-94  1927  8-Sep-97  1969  27-Oct-00
1886  16-Aug-94  1928  5-Oct-97  1970  24-Nov-00
1887  13-Sep-94  1929  1-Nov-97  1971  21-Dec-00

1.6.1 Some Comments about IDL

In using IDL it is very easy to create too many variables to the point where the computer memory capacity is exceeded. It is possible to delete old variables by using the DELVAR command. For example, if you wanted to delete the variable DATA, you would type:

IDL> delvar, data

There is a peculiarity about using the IDL routine FINDFILE and using the Unix symbol ∼ for the user’s home directory. FINDFILE does not return any files when using the command:

IDL> ff = findfile(’∼/*’)

It is recommended to use the full path (for example ‘/2p/morrison/*’) instead of ∼.

It is possible to change the default directory from within IDL by using the IDL command “cd”. Some sample commands are:

IDL> cd, ’∼’ IDL> cd, ’/yd3’

The following will NOT work since it sets the default for the sub-process and then exits that process.

IDL> $ cd /yd3 ;WILL NOT WORK

For Unix machines, it is possible to press control-z to place an IDL session in the background so that you can do some work from the Unix system. When you are ready to go back to the IDL session, type:

% fg

Be careful to remember that you have an IDL session hidden in the background since you can waste memory and corrupt the color table settings if you have multiple IDL sessions going.

1.6.2 Programs, Procedures, and Functions

An IDL main program requires a .RUN command to run, and the code within that file starts executing after successful compilation.

An IDL procedure is a kind of subroutine and has something like

             PRO PROCEDURE NAME, PARAM1, PARAM2

at the top of the file. The variables in the procedure definition can be input or output. It is possible to have keywords with a command like

             PRO PROCEDURE NAME, KEY1=KEY1, KEY2=KEY2

It can be executed with a command like:

IDL> procedure name, a, b IDL> procedure name, a, b, key1=c

An IDL function is another kind of subroutine and has something like

             FUNCTION FUNCTION NAME, PARAM1, PARAM2

at the top of the file. The primary output is passed to result but it is possible to have parameters in the call which can receive output. It is executed with a command like:

IDL> result = function name(a,b)

1.6.3 Running IDL programs

When running on a Unix system, it is important to use lower case when using the .RUN command because Unix is case sensitive. This is not necessary when doing an ‘implied’ compilation when accessing procedures or functions since IDL will convert to lower case for you. You cannot use upper case for a procedure to do an explicit compilation (.RUN PROCEDURE NAME) to recompile a procedure.

2 General Yohkoh Software

2.1 Routines for Reading and Saving Data

2.1.1 RD XDA, RD xxx

The RD xxx routines (RD BDA, RD HDA, RD SDA, RD WDA, RD XDA) allow you to read the reformatted data files. Each routine has the same calling sequence, and there is a generic reading routine called RD XDA which will read data for any of the instruments. After you have established the input file name(s) (infil) and which data sets to extract (dset arr), you can read the data by a command:

IDL> rd sda, infil, dset arr, index, data IDL> rd sda, infil, dset arr, index, data, roadmap IDL> rd xda, infil, -1, index, /nodata

index is a structure which describes the data. There is one index structure for each dataset. The roadmap is for all datasets in the files (not just the selected datasets). It is possible to read all datasets in the file by specifying -1 for dset arr. The /nodata switch tells the routine to only read the index and roadmap, not the data.

2.1.2 RD ROADMAP

RD ROADMAP will read the roadmap from a file (or list of files) by typing:

IDL> rd roadmap,infil,roadmap,ndset

where infil is a string array of file names. The roadmap variable is the same as you get from YODAT

2.1.3 RD QS

RD QS will read the quasi-static portion of a single reformatted file. A sample calling sequence is:

IDL> rd qs, infil, qs

2.1.4 SAVEGEN and RESTGEN

Details on SAVEGEN and RESTGEN are given in the SolarSoft Reference Manual. NOTE: It is recommended to use SAV SDA, SAV BDA, SAV BSC, and SAV HXI whenever possible for standard SXT, BCS, and HXI data sets.

2.1.5 SAV SDA

SDA SDA will store an index and data array in an SDA format file. The data array can be byte, integer*2, integer*4 or real*4. The dimensions of the image do not have to be the original dimensions. A sample is:

IDL> sav sda, outfil, index, data IDL> sav sda, outfil, index, data, qs

where outfil is the name of the file you wish to store the data in. If you wish to append to an existing file, then you can use the /APPEND switch. For example:

IDL> sav sda, outfil, index, data, /append

Use RD SDA to restore the data.

2.1.6 SAV BDA and SAV BSC

SAV BDA and SAV BSC work in the same manner as SAV SDA but for raw BDA data or BSC processed data. Sample calls would be:

IDL> sav bda, outfil, index, data, qs, dp sync IDL> sav bsc, outfil, index, data

2.1.7 SAV HXI

SAV HXI allows you to save HXT synthesized images. A sample call would be:

IDL> sav hxi, outfil, index, data

2.1.8 SXT2FITS

It is possible to take an SXT data file or an index and data that has already been read in, and write a single FITS file for each image. The FITS header has all of the information on the date and time, the filters used, the exposure duration, the resolution, and the DP mode and rate. The following command will create a file for each image in the data array:

IDL> sxt2fits, index, data

The default file names are SF FITSyymmdd.hhmmss and SP FITSyymmdd.hhmmss for FFI and PFI images, respectively. If a single output file name is passed to the SXT2FITS routine, but there are several images to save, then it will append an image number to the end of the file name. In the following example, the input array is 512×512×3, so it will create the files 14 2. GENERAL YOHKOH SOFTWARE ‘flare.0001’, ‘flare.0002’ and ‘flare.0003’. IDL> sxt2fits, index, data, outfil=’flare’ It is possible to specify an input file name in which case all images in that file will have FITS files created. It is also possible to specify a list of the images to be saved. Some examples of these calls are: IDL> sxt2fits, index, data, ss=ss IDL> sxt2fits, infil=infil IDL> sxt2fits, infil=infil, ss=ss

2.2 Routines for Data Selection

2.2.1 TIM2DSET TIM2DSET will take the roadmap and return the dataset number which is closest to that time. Examples, IDL> dset = tim2dset(roadmap, input times) IDL> dset = tim2dset(roadmap, ’23-jun-92 6:00’)

2.3 Routines for Making Time Plots

2.3.1 PLOTY PLOTY plots light curves from all the Yohkoh instruments on the same plot with the same time axis. If you have the observing log on-line and wish to use it to make the plots, you can specify the plot times. It is recommended to plot no more than approximately 24 hours of data at a time. An example is: IDL> ploty, ’8-may-92 15:00’, ’8-may-92 18:40’ If you have identified an instrument file for which you want to see the light curve for all the other instruments, you can use the following command (it assumes that the other instrument data files exist on the same directory as the input file): IDL> ploty, infil=’/yd5/flares/spr911115.2141’ You can read the observing log data, and the use PLOTY with a command like: IDL> rd obs, ’8-may-92’, ’8-may-92 12:00’, bcs, sxtf, sxtp, w h IDL> ploty, bcs, w h, sxtp, w h

2.3.2 OBS PlOT

TODO - needs text

2.4 Routines for Data and Information Extraction

2.4.1 The Concept Behind the GT Routines

A series of GT routines was created so that a single piece of information can be extracted from a structure, whether the structure was a ‘roadmap’ or an ‘index’ (the data are saved in a different location). The routines also allow for conversions to a string mnemonic or to physical units. It is possible to get a list of what the different values are by typing a command like:

IDL> print, gt filta()

It is also possible to convert the output value to the string mnemonic by using the /STRING switch. For example:

IDL> print, gt filta(roadmap, /string)

This command only works for the routines which return a coded value. A list of all of the GT routines currently available are:

GEN  BCS  HXT  SXT  WBS
gt conv2str gt blockid  gt sum h gt adoor  gt grs1
gt day  gt bsc bincal gt sum l  gt center  gt grs2
gt dp mode gt bsc chan gt sum m1 gt comp  gt hxs
gt dp rate gt bsc crate gt sum m2 gt corner  gt rbmsc
gt hxa  gt bsc flux  gt corner cmd gt rbmsd
gt iru  gt bsc time  gt dpe  gt sxs1
gt tfss  gt bsc wave  gt entry  gt sxs2
gt time  gt total cnts  gt expdur

gt explat gt expmode gt filta gt filtb gt mbe gt obsregion gt or expnum gt percentd gt percentover gt pfi ffi gt pix size gt res gt seq num gt seq tab gt shape gt shape cmd gt ssl explab gt sxt axis gt sxt cen gt sxt roll gt temp ccd gt temp hk

2.4.2 GT DP RATE

GT DP RATE will extract the information on the DP telemetry rate. The values are shown below.

1 Low ( 1 Kbits/sec)
2 Med ( 4 Kbits/sec)
4 High (32 Kbits/sec)

IDL> dprate = gt dp rate(index) IDL> dprate = gt dp rate(roadmap, /string)

2.4.3 DPRATE2SEC

DPRATE2SEC will take a structure (or an integer value 1, 2 or 4 for low, medium, or high) and return the number of seconds that passes for a major frame at that telemetry rate.

1 64 sec (Low)
2 16 sec (Med)
4 2 sec (High)

IDL> nsec = dprate2sec(index)

2.4.4 GT DP MODE

GT DP MODE will extract the information on the DP mode. There are several modes, but the most common are:

9 Flare
11 BCS-OUT
12 Night
13 Quiet

IDL> dpmode = gt dp mode(index) IDL> dpmode = gt dp mode(roadmap, /string)

2.4.5 GET INFO

GET INFO will take the roadmap or index structure and return a string describing the main observing mode parameters for each dataset.

IDL> info array = get info(roadmap)

An example for a dataset for SXT data is:

   0 15-NOV-91  22:42:24  FL/H  NaBan/Open  Full Norm C  16   238.0   64x 64

The “0” is the image number in the index or roadmap which was passed to GET INFO, “15-NOV-91 22:42:24” is the start date and time of the exposure, “FL/H” says the DP was in flare mode in high telemetry rate, “NaBan/Open” are the filters, “Full” is the pixel resolution, “Norm” says it was a normal exposure (not a dark frame), “C” is compressed data, “16 238.0” is the DP exposure level (DPE) and the expected exposure duration in milliseconds, and “64x 64” is the image shape.

An example for a dataset for SXT data when using the /LONG option is:

   0 15-NOV-91  22:42:24  FL/H  NaBan/Open  Full Norm C  16   238.0   64x 64  100%  S13W18   0% SAA  2.6

The first 80 columns are the same as above. The “100%” says that all of the image was received from the S/C, “S13W18” are the heliocentric coordinates of the PFI, “0%” of the image was saturated, “SAA” signifies that it was during an South Atlantic Anomaly (SAA) passage, and “2.6” is the number of minutes before spacecraft night.

For HXT,WBS and BCS data, the routine only returns the date, time, DP mode and DP rate.

2.5 Routines for Calibration and Analysis

2.5.1 FSP PROC [*]

FSP PROC is a general routine for doing spectral fits. detector is a string variable containing the detector name, cfile is the input file name, the format for an input file is given in /ys/ucon/soft/mctiernan/spectral data format, ofile is an output file name. tyspec denotes the type of spectrum you want to fit (try doc library, ’fsp proc’ for a complete list). fit pars is the structure containing the results. Optional parameters: sc par is a structure containing spectrometer channel info, ch dta is a structure containing the data for each channel, pfile is the name for an output plot file, sdel is an array of channels you’d like to delete, flux corr is any correction to the overall flux you might want to make, nsigmas tells you how many sigmas above background a “good” channel will be (default is 3.0), noplot suppresses screen plots.

IDL> FSP PROC, detector, cfile, ofile, tyspec, fit pars

2.5.2 FSP 11 [*]

FSP 11 is a general routine for doing spectral fits. inpf is a strarr containing the detector response filenames, tyspec denotes the type of spectrum you want to fit (try doc library, ’fsp 11’ for a complete list). Data must be put in the proper fields of the structures fit pars, sc par and ch dta. fit pars is the structure containing the results, sc par is a structure containing spectrometer channel info, ch dta is a structure containing the data for each channel. Optional parameters: ofile is the name for an output file, pfile is the name for an output plot file, sdel is an array of channels you’d like to delete, flux corr is any correction to the overall flux you might want to make, nsigmas tells you how many sigmas above background a “good” channel will be (default is 3.0), noplot suppresses screen plots.

IDL> FSP 11, detector, inpf, tyspec, fit pars, sc par, ch dta

2.6 Routines for Image Alignment

The routines described here can generally be used on SXT or ground-based images.

2.6.1 SXT PREP

See the description of SXT PREP on page 51 for details on aligning SXT images.

2.6.2 Other Routines

See the SolarSoft Reference Manual.

2.7 Accessing Secondary Databases

2.7.1 CONTACTS

CONTACTS will give information on when the ground station contacts are for a given day. The output from the command below is shown.

IDL> contacts,’2-jun-92’

             Kagoshima Space Center Contacts           Minutes of
                  Starts                   Ends       Day  Ngt  Tot
         JST              (UT)             JST
 2-JUN-92 02:56:14 ( 1-JUN-92 17:56:14)   03:07:14    3.5  7.5 11.0
 2-JUN-92 04:39:59 ( 1-JUN-92 19:39:59)   04:48:59    7.8  1.2  9.0
 2-JUN-92 20:27:14 ( 2-JUN-92 11:27:14)   20:37:14    4.0  6.0 10.0
 2-JUN-92 22:09:14 ( 2-JUN-92 13:09:14)   22:21:14    0.0 12.0 12.0
 2-JUN-92 23:52:29 ( 2-JUN-92 14:52:29)   00:04:14    0.0 11.8 11.8

A blank line is inserted to show that the second set of contacts is for a different series of contacts (they come in clusters of five or six per day). It is possible to get the station contacts times for DSN stations by using the /CANBERRA, /GOLDSTONE or /MADRID switches. It is possible to specify an end time and output files in the following example:

IDL> contacts, ’1-jun-92’, ’10-jun-92’, outfil=’contacts.txt’

2.7.2 PR FEM

PR FEM will print out Yohkoh’s day and night events by typing:

IDL> pr fem, ’1-jan’ 20 2. GENERAL YOHKOH SOFTWARE 2.7.3 PR EVN PR EVN helps you find when there is Yohkoh data available. An event is driven by a change between QUIET and FLARE mode, or when there is a data gap of more than 60 seconds. By typing: IDL> pr evn, ’23-jun-92’ a list of the times that Yohkoh data are available and the number of datasets available for each instrument are listed for 24 hours starting at 23-jun-92 00:00. By typing: IDL> pr evn, ’15-nov-91 20:00’, ’17-nov-91 15:00’, /flare all times that Yohkoh was in FLARE mode between those times is listed. By typing: IDL> pr evn, ’1-jan-92’, ’1-jan-93’, /flare, /counts, mindur=5, outfil=’pr evn.results’ the FLARE modes for 1992 are listed, and since the /COUNTS option was used, the maximum counting rate for certain WBS, HXT, and BCS channels is printed instead of the number of datasets available. It also prints the GOES classification when it is available. See the User’s Guide for a sample output listing. 2.7.4 PLOT EVN PLOT EVN will plot a very basic time line showing when the station contacts are available, when the SAA passages are, when the S/C day and nights are, and the periods when there is Yohkoh data available. It shows when there is FLARE and QUIET data. 2.7.5 RD FEM To read the FEM, EVN ,GEV or NAR structures, use a command similar to: IDL> rd fem, sttim, entim, fem IDL> rd evn, index(0), index(n-1), evn IDL> rd nar, ’1-nov-91’, ’2-nov-91’, evn Look at the File Control Document for a description of the structures which are returned. 2.7.6 RD OBS A sample standard calling sequence for reading the observing log data files for all instruments is: IDL> rd obs, ’8-may-92’, ’8-may-92 12:00’, bcs, sxtf, sxtp, w h, fid 2.8 Reading Exabyte Archive Tapes 21 It is possible to only read SXT full-frame data by using the one of the following command: IDL> rd obs, ’8-may-92’, ’8-may-92 12:00’, bcs, sxtf, sxtp, w h, fid, /sxtf IDL> rd obs, ’8-may-92’, ’8-may-92 12:00’, bcs, sxtf, /sxtf IDL> rd obs, ’8-may-92’, ’8-may-92 12:00’, bcs, sxtf, /nobcs To read the WBS data, you can use the command: IDL> rd obs, ’8-may-92’, ’8-may-92 12:00’, bcs, sxtf, sxtp, w h, /w h The observing log output structures are positional and are sorted alphabetically. The observing log files must be in the $DIR GEN OBD, $DIR GEN OSF, $DIR GEN OSP, and $DIR GEN OWH directories appropriately.

2.8 Reading Exabyte Archive Tapes

See the description of GO RDTAP in the User’s Guide and the description of RDTAP on page 73 in the appendix of the Reference Guide.

3 Bragg Crystal Spectrometer (BCS)

3.1 Routines for Data Selection

3.1.1 LIST BDA

A sample call would be:

IDL> list bda, roadmap, start, nrec, ss=ss

where roadmap is the input.

3.1.2 SELECT BDA

Use SELECT BDA to plot the roadmap light curve of a selected channel and select the time period of interest using the cursors. (Note: there is an option to read the data in with SELECT BDA if required). A sample call would be:

IDL> .run select bda

where roadmap is the input.

3.2 Routines for Making Time Plots

3.2.1 BCS 24HR PLOT

If you interested only in whether the BCS has seen anything, use BCS 24HR PLOT. This works using only roadmap data and allows intervals to de selected by date (all files covering this data are then used). By default, channel 3 (Ca XIX) is plotted. Note: YODAT does NOT need to have been run before BCS 24HR PLOT. In its default mode, BCS 24HR PLOT uses the observing log files. Samples are:

IDL> bcs 24hr plot, ’15-nov-91’ IDL> bcs 24hr plot, ’15-nov-91’, chan=4

3.2.2 PLOTT BDA

PLOTT BDA will make a light curve plot using either the index or roadmap. The second ‘T’ signifies that it is a time plotting routine. Four plots are made on the page, one for each channel. Some sample calls are:

IDL> plott bda, index IDL> plott bda, roadmap IDL> plott bda, roadmap(100:200), psym=10

3.2.3 LCBDA [*] LCBDA will make a light curve plot using either the index or roadmap. It is almost identical to PLOTT BDA except that you can specify a channel and only get that channel plotted. Some sample calls are:

IDL> lcbda, index IDL> lcbda, roadmap, chan=1

3.3 Routines for Making Spectral Plots

3.3.1 PLOTS BDA PLOTS BDA allows for the spectra for all four channels to be plotted to one page. A sample calling sequence is:

IDL> plots bda, index, data

3.3.2 PLOTBDA The light curve and spectra from the BDA file may be plotted using PLOTBDA. This is an interactive program and it will read its own data.

IDL> .run plotbda

3.3.3 BCS MULTI Many spectra may be plotted on a page with BCS MULTI.

IDL> bcs multi, index, data IDL> bcs multi, index, data, chan=1

3.3.4 BCS CONT The evolution of spectra against time can be displayed as a contour map using the routine BCS CONT. Note: The time axis of BCS CONT is uniform, but those of DISP BDA and GS are not.

IDL> bcs cont, index, data IDL> bcs cont, index, data, chan=1

3.4 Routines for 2-D Spectral Display

3.4.1 DISP BDA

The evolution of spectra against time can be displayed as a pseudo-image using the routine DISP BDA. Note: The time axis of DISP BDA is not uniform.

IDL> disp bda, index, data

3.4.2 BCS SPMOVIE

After selecting a time interval, a movie of the changing spectra for a given channel may be displayed by BCS SPMOVIE. Note: This program will only run on an X-windows workstation.

IDL> bcs spmovie, index, data IDL> bcs spmovie, index, data, chan=1

3.4.3 GS [*]

If you are interested in more detail of what spectra the BCS has observed, use GS — this routine must be run after the data has been read in with YODAT. There is an upper limit on how much data can be handled at a time (about 900 spectra), and some selection of the required time interval may be needed using the methods described above, but for the per-orbit files, this limit may not be a problem. The advantage of GS is that you can get a good idea of what has been seen in the spectra — a particularly useful tool if you are trawling for data. Warning: The time axis on the greyscale plot is not uniform!

IDL> .run gs

All the routines that run from GS require that spectra be selected first using the cursor routine.

IDL> .run gs cur

3.5 Routines for Data and Information Extraction

3.5.1 EXT BCSCHAN

The spectra for a single channel may be extracted with EXT BCSCHAN

IDL> data out = ext bcschan(index, data, chan)

3.5.2 LIST BDA

If you are interested in what modes the BCS was executing, or what the count rate in a particular channel has observed at a particular time, this can be determined by using LIST BDA. By default the count rate for channel 3 (Ca XIX) is given in the listing. Sample calls are:

IDL> list bda, roadmap, start, nrec IDL> list bda, roadmap, start, nrec, chan=4

where roadmap is the input.

3.5.3 GT TOTAL CNTS

GT TOTAL CNTS extracts the information from the TOTAL CNTS field which used the actual spectra to calculate the total number of counts. The routine also normalizes so that it returns counts/sec. In the first example below, all four channels are returned, so the output is 4xN. In the second example, only channel 1 is extracted. It is possible to get a string defining the channel selected by using the title keyword option.

IDL> x = gt total cnts(roadmap) IDL> x = gt total cnts(roadmap,1) IDL> x = gt total cnts(index,2,title=title)

3.5.4 GT BLOCKID

GT BLOCKID returns information on how the BCS data were blocked.

IDL> x = gt blockid(roadmap)

0 Normal Queue Data Block
1 Fast Queue Data Block
2 Micro Dump Block (fixed extraction)
3 Cal Data Block (fixed extraction)
4 Queue data where the modeID in the header is not recognized
5 Normal or fast queue data which have fill data (garbage)

3.6 Routines for Calibration and Analysis

3.6.1 BCS DECOMP The BCS data are normally compressed from a 12-bit word to an 8-bit word. BCS DECOMP takes the compressed data array (data) and creates a floating point array (ndat) of the decompressed numbers by typing:

IDL> ndat = bcs decomp(data)

3.6.2 BCS NORM BCS NORM normalizes data recorded by the BCS for integration time, and extracts the Fast Queue data if present. A sample call is:

IDL> dspec = bcs norm(index, data)

3.6.3 SUM BDA All spectra with mode ID modeid will be summed by SUM BDA with the following command:

IDL> data out = sum bda(index, data, modeid, nsum)

3.6.4 PLOT REF [*] Overplot several spectra to show evolution of blue wing and line width.

IDL> plot ref, index, data, channel, dset arr

3.6.5 BCS DTFAC The dead-time correction factors can be calculated with BCS DTFAC:

IDL> dd = bcs dtfac(index,dp sync,type,stime,etime)

3.6.6 BCS POINT The position of a source on the solar disk may be determined from the bin number of the resonance line with BCS POINT. If the epoch is provided, this routine will also return the heliocentric latitude:

IDL> dd = bcs point(wbin [,chan=n [,date=date, etc.]])

3.6.7 BCS SYNTHETIC

BCS SYNTHETIC is a simple main program to compute BCS synthetic spectra in flux units. The program will prompt the user for the BCS channel number (1,2,3,4 = Fe XXVI, Fe XXV, Ca XIX, S XV), the electron temperature (MK) and the Doppler temperature (MK). To run the program type:

IDL> .run bcs synthetic

The program calls the procedure BCS SPEC to perform the spectral calculations. It defaults to a wavelength coverage which is appropriate for mid-latitude observations. The BCS SPEC defaults for atomic code, ionization balance and abundance will be used. The computation is made for an emission measure of EM = 10^{50} cm^{−3}.

3.6.8 BCS SPEC

BCS SPEC is a procedure to compute a BCS synthetic spectrum in flux units. Sample calling sequences are:

IDL> flux = bcs spec(Chan,Te6) ; Use Td6=Te6 and default wavelengths IDL> flux = bcs spec(Chan,Te6,Td6,Wave,Waveo=Waveo,Ainfo=Ainfo,Vel=Vel)

where the inputs are Chan (the BCS channel number of a BSC structure), Te6 and Td6 (electron and “Doppler” temperatures in MK), and Wave is a vector of the edges of the BCS wavelength bins. Returned will be flux (assuming EM = 10^{50} cm^{−3}), Waveo (wavelengths at the centers of the bins), and Ainfo (a structure variable containing information about the calculated spectrum). Vel is an optional input keyword to specify the bulk velocity in km/s. More information is available in the documentation header. BCS SPEC is called by BCS SYNTHETIC and FIT BSC.

Three routines control the defaults for the atomic data used in the calculation. These are GET ELEMABUN (abundances), GET IONBAL (ionization balance) and GET ATOMIC (the atomic data for the line calculation).

3.7 BSC Routines

3.7.1 HELP BSC

To get quick on-line help for various BSC procedures and functions, type:

IDL> help bsc 3.7.2 MK BSC

MK BSC makes BSC INDEX and BSC DATA structures from the BDA INDEX and DATA. These structures contain instrumentally corrected BCS data for selected channels and accumulation times. A typical run would look like:

IDL> mk bsc,index,data,bsc index,bsc data,dp sync=dp sync

To analyze BCS data, you would begin by using YODAT to read the BDA data file for the dates and times you are interested in. Yodat will return a BDA index structure INDEX, a BDA data array DATA, and a structure DP SYNC which contains the DP synchronous data needed for deadtime corrections.

MK BSC will prompt for various inputs such as start and stop times, channels to process, integration times, and which corrections (deadtime, crystal curvature, flux calibration) to apply. The DP SYNC keyword is optional, but must be entered the first time if deadtime corrections are required. Subsequent calls to MK BSC will use the last values of DP SYNC data passed via this keyword.

The following keywords are also accepted:

/LAST  – run using the information stored in the latest BSC INDEX.
/DEF  – run using the default settings.
/WID  – run via a widget interface.
/BSA  – run to make BSA index only (do not make BSC DATA).
/SAVE  – save BSC (or BSA) index (and data) in a file.

(The file will be named BSCyymmdd.hhmm, based on the time of the first data interval).

/VERBOSE – echo extra information about SAA/NIGHT or MODEID/BLOCKID
BDAFILE – input BDAFILE to read.

The instrumentally corrected data will be output to a new set of structures BSC INDEX and BSC DATA. The data in BSC DATA are stored as 256 element floating point arrays in fields with the following definitions:

.COUNTS – accumulated counts (per detector bin)
.BIN  – physical detector position
.ERROR  – spectral uncertainties (in counts)
.WAVE  – nominal wavelength array (Å)
.FLUX  – flux-calibrated data (ph cm^{−2} s^{−1} Å^{−1)}
.WAVE FIT – wavelength array for the fitted spectra (Å)
.FLUX FIT – fitted spectra

Not all the above fields will be present. The fields that are present will depend on the corrections requested in MK BSC. The actual data in each field array will occupy the first 0:NBINS-1 locations, where NBINS is the number of actual detector bins containing valid data.

3.7 BSC Routines 29 The value of NBINS is stored in BSC INDEX.BSC.NBINS. 3.7.3 SEL BSC SEL BSC extracts wavelength and flux fields from BSC DATA. IDL> sel bsc,bsc index,bsc data,new index,new data,wave,flux,eflux The selected index and data structures are returned in NEW INDEX and NEW DATA. The wavelength, flux, and uncertainties in the flux are returned in the arrays WAVE, FLUX, and EFLUX.

The following keywords are also accepted:

CHAN (in)  - BCS channel (def is all)
TSTART (in)  - start time to extract (UTPLOT format)
TEND (in)  - end time to extract
BIN (out)  - Physical detector positions
COUNTS (out) - observed counts (per bin) in the accumulation interval
SS (out)  - indices matching requested channel
FWAVE (out)  - fitted wavelength array
FIT (out)  - fitted spectrum array
SFIT (out)  - fitted second component array

3.7.4 LIST BSC LIST BSC lists times, total counts, and accumulation times of BSC spectra. IDL> list bsc,bsc index,chan where CHAN is the desired channel to list. 3.7.5 SAV BSC To save BSC index and data to a BSC file, use: IDL> sav bsc,’bsc file name’,bsc index,bsc data Note that MK BSC will also save the data if the /FILE switch is set. 3.7.6 RD BSC To read a BSC file, use:

30 3. BRAGG CRYSTAL SPECTROMETER (BCS) IDL> rd bsc,’bsc file name’,ss,bsc index,bsc data This procedure will return the BSC index and data structures. SS is a vector with indices of a particular subset of data within the BSC file. For example, using SS=INDGEN(10) will extract the first 10 data sets from the BSC file. To extract ALL the data sets, use SS = −1. 3.7.7 BSC2BSD To convert a BSC file to a BSD file, use: IDL> bsc2bsd,’bsc file name’,’bsd file name’ The BSD file name is optional, and will be created from the BSC file name if not supplied. 3.7.8 MAP BSC MAP BSC plots a pseudo-map of BSC spectra by stacking successive spectra. IDL> map bsc,bsc index,bsc data The following keywords are accepted:

CHAN  BCS channel to plot.
XSPACE the value by which to shift successive spectra in the x-direction.

[def=0]

YSPACE the value by which to shift successive spectra in the y-direction.

[def=0]

TSTART the time of the first spectrum to plot (with date optional).

[sample format,TSTART=’9:32’, def = first spectrum in BSC DATA]

TEND  the time of the last spectrum to plot.

[def = last spectrum in BSC DATA]

3.7.9 LC BSC LC BSC plots light curves of selected spectral regions. IDL> lc bsc, bsc index, bsc data, t, y, chan=chan,/flux This program first plots a spectrum at the peak of the flare. The user can select a spectral region with the cursor. The light curve derived from the sum of wavelength points within this region is returned in the array Y. The time (UT) is returned in T. If the BSC data are flux calibrated, then using the FLUX keyword will return the light curve in units of photons cm^{−2} s^{−1}. The optional keyword over will overplot successive light curves.

3.7.10 WBDA

WBDA provides a simple widget interface that allows browsing of BDA files. It is invoked by:

IDL> wbda

The program will provide a set of menus from which BDA files can be selected either by date, filename, or directory location. WBDA will plot light curves and spectra for a specific channel, accumulate spectra for selected integration times, and apply dead-time, curvature, and flux corrections. The BDA INDEX and DATA variables produced by YODAT can be also input directly to WBDA:

IDL> wbda,index,data

3.7.11 GT BSC BINCAL

This function returns a structure containing miscellaneous calibration information for specified BCS channels: For example, for channel 1,

IDL> bincal=gt bsc bincal(1)

will return the following structure:

.CHAN  - channel number
.NBINS  - number of bins per channel
.MODEID - mode id
.WAVE  - nominal wavelength array (Å)
.PHYSPOS - physical detector position array
.FLUXFAC - conversion factor array for counts/bin to photons/cm/Å
.SENSIT  - mean sensitivity array for grouped bins
.DBIN  - detector bin width
.W0  - start wavelength of spectrum
.DW  - dispersion (angstrom/bin)
.BINARC  - conversion factor for arcmin to bin (bins/arcmin)
.GAUSSW - FWHM Gaussian detector width (Å)
.RWID  - FWHM rocking width (Å)
.EFAREA  - effective area (cm^{2})
.NSTART  - first valid bin
.NEND  - last valid bin
.VERSION - version number of BCS calibration file used

3.7.12 FIT BSC

FIT BSC fits synthetic spectra to observed spectra in BSC DATA:

IDL> fit bsc,bsc index,bsc data,fit index,fit data,chan=chan

In the above call, the fitted results are returned in the new structures FIT INDEX and FIT DATA for the channel specified by the keyword CHAN. If CHAN is not specified, then all channels will be fit (not recommended).

The FIT DATA structure contains two additional fields:

.FLUX FIT a 256-element floating point array containing the fitted model spectrum matching the observed spectrum in the .FLUX field

.WAVE FIT a 256-element floating point array containing computed wavelengths of the fitted model spectrum in .FLUX FIT.

.FLUX FIT2 a 256-element floating point array containing the fitted model spectrum corresponding to a secondary velocity shifted spectrum in the .FLUX field

The FIT INDEX structure contains an additional .FIT field that contains important fit information (parameters fit, number of iterations, chi squared, etc).

FIT BSC accepts the following optional keywords:

SS index selection vector or scalar (e.g. SS=[45,46] to fit only spectra 45

and 46)

TSTART UT string (e.g. ’9:45’). Only spectra at or after TSTART will be fit.
TEND  Only spectra at or before TEND will be fit.

LRANGE 2-element wavelength vector range (Å) within which to limit line fitting

[Default is entire range]

CRANGE 2-element wavelength vector range (Å) within which to limit continuum

fitting [Default is entire range]

TE6 input starting guess for electron temperature (MK) [Default is to use

internal values]

TD6 input starting guess for resonance line Doppler broadening (MK) [Default is to infer from observed width] ATTEMPT level of fit attempt

1. solve only for primary emission measure (EM50) and line to continuum normalization factor (CNORM). Electron temperature and (TE6) and total Doppler broadening (TD6) fixed.

2. solve for TE6, TD6, EM50, CNORM.

3. solve for primary+secondary shifted component.

/BLUE  equivalent to attempt=3

/NEWGUES set to NOT use results of previous fit as starting input for fit to next

spectrum

/PLOT  set to plot progress of fits
/VERB  set for more information about fitting progress

/LAST set to use last converged parameters as starting guess for fitting next

spectrum.

For example, to fit a Ca XIX spectrum at UT=’09:04’ use:

IDL> fit bsc,index,data,findex,fdata,chan=3,tstart=’09:04’

To fit a blueshifted secondary component:

IDL> fit bsc,index,data,findex,fdata,chan=3,tstart=’09:04’,/blue

3.7.13 BCS BROAD

The relation between BCS line broadening and velocity can be computed with BCS BROAD. Some example calling sequences are:

IDL> Td6 = bcs broad(Chan,TDoppw=TDoppw) ; Return temperature in (MK) IDL> Vel = bcs broad(Chan,Td6=Td6,/vel) ; Return velocity (km/s) IDL> TDoppw = bcs broad(Chan,Td6=Td6,/Gauss) ; Return Doppler FWHM (Ang)

where T D6 is the equivalent Doppler Temperature (MK) and CHAN is the BCS channel number. The parameter T DOP P W does not include the contribution due to the crystal (Lorentz) broadening, but is included by BCS BROAD when it makes its calculations, based on the value of CHAN. There are several additional options available and there is a keyword to return the crystal rocking curve width. Most conversions between BCS line width and velocity or between line width and temperature can probably be accomplished using this routine.

3.7.14 PLOT BSC

The results of MK BSC and FIT BSC can be examined with PLOT BSC.

IDL> plot bsc,bsc index,bsc data

PLOT BSC will plot light curves and spectra for user selected channels and times, as well as fitted spectra from FIT BSC.

Optional keywords:

CHAN  - channel to plot
SS  - selected indicies of spectra to plot (default is all)
/NOWID - to not use widget mode (def is widgets, if more than 1 spectrum)

The following keywords apply only when /NOWID is set:

/EBAR - oplot spectral uncertainties
/HC  - make a hard copy (single spectrum only)
/PS  - make a PostScript file (single spectrum)

The following keywords apply only when /NOWID is set, and BSC INDEX and DATA contain fitted spectra:

/BLUE - to oplot second blueshift component (if fitted by FIT BSC)
/VUNIT - to print fitted Doppler widths in km/s units (default is MK)
/NOFIT - inhibit oplot of fitted spectra (default is oplot fit)
/NOOBS - inhibit plot of observed spectra (default is plot obs)

3.7.15 FIT BSC PLOT

FIT BSC PLOT produces time series plot results from FIT BSC. The routine will plot total counts, electron temperature, emission measure, and turbulent velocity as a function of time. Example calling statements are:

IDL> fit bsc plot,bsc index IDL> fit bsc plot,bsc index,tstart=tstart IDL> fit bsc plot,bsc index,/hc ; Produce hardcopy output

Consult the document header for additional options.

3.7.16 MK FIT BSC EPS

MK FIT BSC EPS will produce a BCS spectrum in portrait mode. It optionally will specify the output file to be an EPS (encapsulated PostScript file). Some sample calls are given below:

IDL> mk fit bsc eps, index, data, /hc ; Send plot to the printer IDL> mk fit bsc eps, index, data, /ps ; Make file but don’t print IDL> mk fit bsc eps, index, data, /eps ; Create an EPS file IDL> mk fit bsc eps, index, data, /noatomic ; Suppress atomic code label

MK FIT BSC EPS sets up the parameters to call BCS SPEC PLOT correctly.

The default X and Y size is about right for an A4 page. The optional xoffs=xoffs and yoffs=yoffs enable the user to position the plot at the desired location on the page.

4 Hard X-Ray Telescope (HXT)

4.1 Routines for Making Time Plots

4.1.1 PLOTT HDA PLOTT HDA will make a light curve plot of all four channels using either the index or roadmap. The second ‘T’ signifies that it is a time plotting routine. Four plots are made on the page, one for each channel. Some sample calls are:

IDL> plott hda, index IDL> plott hda, roadmap IDL> plott hda, roadmap(100:200), psym=10

4.2 Routines for 2-D Spectral Display

4.2.1 DISP HDA The evolution of sensor intensity against time can be displayed as a pseudo-image using the routine DISP HDA. Note: The time axis of DISP HDA is not uniform.

IDL> disp hda, index, data

4.3 Routines for Data and Information Extraction

4.3.1 GT SUM L These routines will extract the average counts/sec of all 64 sensors. The input can be roadmap, index, or observing log and it will get the proper structure tag and decompress it properly to return counts/sec/sensor. It is possible to get a string defining the channel selected by using the title keyword option.

IDL> y = gt sum l(roadmap) IDL> y = gt sum l(w h, title=title)

4.3.2 GET INFO See the description in the “General Yohkoh Software” chapter on page 17.

4.4 Routines for Calibration and Analysis

4.4.1 HXT QLOOK HXT QLOOK makes a series of HXT quick-look images (i.e. a quick-look movie). This program is much faster than the MEM code, but should be used only for quick-look movies. It is very useful for verifying positions etc. before running a long MEM job.

IDL> hxt qlook, index, data, movie, hxi index IDL> hxt qlook, index, data, movie, hxi index, patterns

index  HXT index from an HDA file
data  HXT data from an HDA file

patterns the sub-collimator modulation patterns. Computed if not supplied. The computation is slow, so if you plan to make a series of movies, use patterns to receive the computed patterns the first time you call HXT QLOOK and then use patterns to pass the modulation patterns back to the routine on subsequent calls to HXT QLOOK. You must recompute the modulation patterns whenever you use a different pointing.

movie  a data cube of the HXT images (output)

hxi index an index describing various parameters of the images. The time in the index is the start of the accumulation period (output).

/new mod patterns Force a new computation of the modulation patterns, even if the patterns parameter is set. Normally, if patterns is set, the modulation patterns are not computed. You must recompute the modulation patterns whenever you use a different pointing.

/show images  show each image as it is computed
/quiet  don’t print any diagnostics
/verbose  print lots of diagnostics

smooth smoothing parameter which determines how much the inverse computation of the image is smoothed. Default = 0.1.

4.4.2 HXTPRO

See the User’s Guide for a description of this program.

4.4.3 HXT IMG

See the User’s Guide for a description of this program.

4.4.4 GET HXT POS

The routine GET HXT POS will return the HXT address of the flare location. It will perform the following checks in this order until it finds a match.

1. Read the $DIR HXT CAL/hxt flare pos.txt file to see if the time is in that list

2. Use the GOES heliocentric location to determine the address

3. Use an SXT image index to define the address

4. Pass in heliocentric coordinates and convert them

5. Ask for user to specify HXT address directly

Sample calls are:

IDL> xy0 = get hxt pos(’15-Nov-91 22:37’) IDL> xy0 = get hxt pos(index(20), helio=’N32W23’) IDL> xy0 = get hxt pos(index(20), status=status)

The optional keywords /get sxt, /get goes and helio will force the calculation of the function to use that technique. The keyword option /nouser should be used when running in batch mode as it will inhibit asking the user for input. The status output is a −1 if no match has been found, 1 for HXT flare list, 2 for SXT image, 3 for GOES event list, 4 for heliocentric coordinate passed in, and 5 for user input.

4.4.5 HXT FSP and HXTBOX FSP [*]

HXT FSP is the same as HXS FSP, but for HXT spectra. See HXS FSP for a better description. Index and data come from an hda file.

IDL> HXT FSP, index, data, fit pars

HXTBOX FSP fits spectra to a given box, chosen by LCUR IMAGE, for a set of HXT images in the HXI structure; index and data come from and hxi file of deconvolved images.

IDL> HXTBOX FSP, index, data, fit pars

4.4.6 SXTHXT FSP and SXTHXTBOX FSP[*]

SXTHXT FSP fits combination SXT and HXT spectra, no spatial resolution! index and data are uncompressed SXT images, hindex and hdata are uncompressed HXT data, from an HDA file.

IDL> SXTHXT FSP, index, data, hindex, hdata, fit pars

SXTHXTBOX FSP fits SXT-HXT spectra for a box chosen using LCUR IMAGE. hindex must be the HXI structure, and hdata are HXT images. You should be sure of alignment beforehand.

4.4 Routines for Calibration and Analysis 39 IDL> SXTHXTBOX FSP, index, data, hindex, hdata, fit pars 4.4.7 HXT DECOMP HXT DECOMP will decompress the HXT data. A sample call is IDL> odata = hxt decomp(data)

5 Soft X-Ray Telescope (SXT)

5.1 Routines for Data Selection

5.1.1 SSWHERE

SSWHERE is a widget-driven program that creates an array of subscripts (ss) that fulfill a set of criteria that you define to select SXT images. For example to make an ss array for YODAT where you want to select images from a data set represented by roadmap, you merely type:

IDL> ss=sswhere(roadmap)

SSWHERE works on both the roadmap and index structures. Available selection criteria are:

• completeness of the image

• SXT mode (normal/dark image/calibration)

• pixel resolution

• compression mode

• filters

• exposure DPE

• DP mode

• field of view location (see lower buttons)

• time period (see lower buttons)

If you get a “No matching Data” response, check that you have selected all possible options.

5.1.2 SHOW OBS3

SHOW OBS3 plots a time line summary of the SXT images that are in the index or roadmap. In addition to plotting a tick for each image available, it shows the DP mode, DP rate, where the S/C days are, and where SAA passages occur. Some sample calls to SHOW OBS3 are:

IDL> show obs3, roadmap IDL> show obs3, index

It is possible to select images using SHOW OBS3 by clicking on diagonally opposite corners of a box surrounding the images you wish to select. The calling sequence for that is:

IDL> show obs3, roadmap, sel, ss

where sel is a returned logical array the same length as roadmap, and is set true for all dsets selected, and ss is the returned subscripts of the selected datasets. A few examples on how to use the sel variable are:

IDL> ss = where(sel) IDL> ss = where(sel and (gt dp mode(roadmap) eq 9))

5.1.3 SHOW OBS4

The calling sequence of SHOW OBS4 is the same as SHOW OBS3, except there are a few additional options in SHOW OBS4. Run SHOW OBS4 with one of the following commands:

IDL> show obs4, roadmap IDL> show obs4, index

It is possible to click on a time range for a closer look by:

1. Clicking with the left button at the start time

2. Clicking with the right button at the end time

3. Clicking with the left button on the box in the lower right

4. If you wish to exit, click with the middle button somewhere on the plot

It is possible to select several regions using the following steps.

1. Click with the right button on the box in the lower left (Select Option). A square box should appear on the screen.

2. Click and hold the left button on the lower left corner of the box and drag the box to the lower left corner of the region you want to select, then release.

3. Click and hold the middle button on the upper right corner of the box and stretch the box to encompass the images you want to select.

4. Repeat steps 2 and 3 until you have positioned the box where you want it.

5. Click on the right button to exit the selection option when you have successfully positioned the box.

6. Now you can exit SHOW OBS4 by clicking on the middle button, or you can select another set of images by running steps 1 through 5 again.

5.1.4 PLOT FOV

It is possible to make a simple display of the sun and to overlay the box showing the field of view size and location by using PLOT FOV. The index, roadmap or observing log entry can be used with this routine. A sample call is:

IDL> plot fov, index, /box

5.2 Routines for Making Time Plots

5.2.1 LCUR IMAGE

See the description in the “General Yohkoh Software” chapter on page ?? or the description in the User’s Guide.

5.3 Routines for Image Display and Enhancement

5.3.1 STEPPER

See the description in the “General Yohkoh Software” chapter on page ?? or the description in the User’s Guide.

5.3.2 XY RASTER

XY RASTER makes a mosaic of several images and prints the times in the corner of each image. Sample calls are:

IDL> xy raster, index, data IDL> xy raster, index, data, factor

where factor is the rebin factor (a value of 1 will perform no rebinning). If there are too many images, the program will display the number of images which will fit on a page, and will prompt the user for when to move to the next page. The keyword /notimes will inhibit the time labels, /noscale will inhibit image scaling (it will use TV instead of TV SCL), /individual scale will scale each image individually.

5.3.3 DISP MONTH

The routine DISP MONTH will display a month of SXT images in a calendar form (assuming you have the SFM database on-line). A sample call is:

IDL> disp month, ’oct-93’

5.3.4 UNSHARP MASK

UNSHARP MASK is a basic image enhancement routine which uses the unsharp mask method. The calling sequence is:

IDL> dataout=unsharp mask(index, data)

where data and dataout are 2D arrays, and index is the index entry that corresponds to data. If no keywords are given the defaults are used.

Optional Keyword Parameters:

smooth  Size of smoothing box, integer.
lowdelt Lower δ cutoff, a percentage of minimum, decimal 0.xx.
updelt  Upper δ cutoff, a percentage of maximum, decimal 0.xx.

lowint Saturation cutoff, don’t add back when signal is BELOW

a given percentage of the saturation value, decimal 0.xx.

upint Saturation cutoff, don’t add back when signal is ABOVE

a given percentage of the saturation value, decimal 0.xx.

deltcoef Multiplier for δ field, float x.x.

The keyword parameters allow you to customize the enhancer, though they are all assigned default values if omitted in the call. These defaults depend on the images resolution, and may be changed as we gain experience.

5.3.5 DE SPIKER

The DE SPIKER function returns a cleaned image by removing spikes at a specified input level. The technique used is to difference a smoothed image and the input image and find all pixels that deviate by more than a given threshold.

IDL> clean img = de spiker(image, thresh) IDL> clean img = de spiker(image, thresh, /neg) IDL> clean img = de spiker(image, thresh, indices=indices)

where, image is the input data array, thresh is the specified spike amplitude to remove. Optional input/output parameter indices contains the indices of spikes at the level specified, and optional input switch /neg requests that “dark holes” be removed.

5.3.6 SXT GRID

Draws heliocentric grid over an image. It reads the ATT file to get information on the spacecraft pointing, and also calls GET RB0P to get the solar radius. A sample call could be:

IDL> tvscl, data(*,*,0) IDL> sxt grid, index(0)

The keyword option /read will allow the user to move the cursor around the image and to display the heliocentric coordinates. If /angle is used with /read, then the value displayed is the angle relative to the sun center, which is useful to be used with SXT PREP.

44 5. SOFT X-RAY TELESCOPE (SXT)

5.4 Routines for Data and Information Extraction

5.4.1 GET INFO See the description in the “General Yohkoh Software” chapter on page 17 5.4.2 EXT SUBSET [*] EXT SUBSET extracts a subset (in x and y) of a data cube and creates a new data and index array (data out and index out) by typing: IDL> ext subset,index in,data in,index out,data out It is a mouse driven program. 5.4.3 GT FILTA GT FILTA uses an index or roadmap structure to determine the setting of the A filter: IDL> a = gt filta(index) The filter setting will be returned as an integer (1-6) where

1 Open Op Op
2 Optical - narrow band  NaBan NB
3 Quartz defocusing lens  Quart Qz
4 Diffuser Diffu Df
5 Optical - Wide band  WdBan WB
6 8% neutral density X-ray filter NuDen ND

The last two columns show what GT FILTA returns when using the /string and /short options. The routine GET INFO uses the /string option. It is possible to get a string type description of the filters used by typing: IDL> b = gt filta(roadmap, /string) You can get a listing of these at any time by typing the following: IDL> print,gt filta() 5.4.4 GT FILTB GT FILTB works exactly like GT FILTA but works for the X-ray filters

1 Open  Open Op
2 Thin Aluminum (1265 Å) Al.1 A1
3 Dagwood Sandwich  AlMg AM
4 Beryllium (119 μm)  Be119 Be
5 Thick Aluminum (11.5 μm) Al12 A2
6 Magnesium (2.5 μm)  Mg3 Mg

The last two columns show what GT FILTB returns when using the /string and /short options. The routine GET INFO uses the /string option.

5.4.5 GT PFI FFI

GT PFR FFI extracts the information on whether the image is a partial frame image (PFI) or a full frame image (FFI). Even values are PFI, odd values are FFI in general.

0 PFI
1 FFI
2 PFI Observing Region
3 FFI Patrol Image

The /ffi keyword option can be used, and it will return a 1 if the image is an FFI and return a 0 if it is a PFI (PFI strip or observing region). The /true ffi keyword option will return a 1 if the image is an FFI and return a 0 if it is a PFI (PFI strip or observing region). The /true ffi option will also check if it is an FFI strip or if it is a PFI which was extracted from an FFI.

5.4.6 GT RES

GT RES extracts what resolution each SXT image was taken at by typing:

IDL> res = gt res(roadmap)

It outputs the resolution as a integer where:

0 full resolution
1 half resolution
2 quarter resolution

It is possible to get a string type description of the resolution by typing:

IDL> res = gt res(roadmap, /string)

46 5. SOFT X-RAY TELESCOPE (SXT) 5.4.7 GT COMP GT COMP extracts what compression was used for each SXT image was taken at by typing: IDL> comp = gt comp(roadmap) It outputs the compression as a integer where:

0 Compressed (see routine SXT DECOMP)
1 Low 8 bits (of 12 bits)
2 High 8 bits (of 12 bits)

5.4.8 GT DPE GT DPE extracts what DP exposure (DPE) level was used. The DPE is a function of the shutter exposure level (MBE) and the neutral density filter use. It is possible to return the effective exposure duration in milliseconds by using the /conv switch). IDL> dpe = gt dpe(index) IDL> msec = gt dpe(roadmap, /conv) The effective exposure for each of the DPE values are shown in the following table. For DPE values above 22, the effective exposure is

                exposure = (DPE-21) * 0.25 seconds

DPE  Effective MBE  Mask DPE  Effective MBE Mask
Exposure  Exposure
0  0.08 msec  0 with ND  19 668.00 msec  12
1  0.23  1 with ND  20 948.00  13
2  0.96  0  21 1.338 sec  14
3  1.37  2 with ND  22  1.888  15
4  2.88  1  23  2.668  16
5  3.06  4 with ND  24  3.778  17
6  4.67  5 with ND  25  5.338  18
7  6.28  6 with ND  26  7.548  19
8  9.50  7 with ND  27 10.678  20
9 17.00  2  28 15.108  21
10 28.00  3  29 21.358  22
11 38.00  4  30 30.208  23
12 58.00  5  31 42.718  24
13 78.00  6  32 60.418  25
14 118.00  7  33 85.438  26
15 168.00  8  34 120.828  27
16 238.00  9  35 170.878  28
17 338.00  10  36 241.668  29
18 468.00  11

5.4.9 GT MBE

GT MBE extracts what mail box exposure (MBE) level was used. This is the actual shutter exposure level commanded. It is possible to convert to commanded shutter duration in milliseconds by using the /conv switch).

IDL> mbe = gt mbe(index) IDL> shutdur = gt mbe(roadmap, /conv)

5.4.10 GT EXPDUR

GT EXPDUR will calculate the effective exposure duration (not the commanded DPE) of a series of images given an index structure to work with. You can get an array of exposures (exps) in msec by typing one of the following. In the second example, it returns the shutter duration (which is also the CCD integration period). For dark images the value returned is the number of milliseconds the CCD integrated for which is in multiples of 8 msec (8 msec is the minimum for dark current images).

IDL> expos = gt expdur(index) IDL> expos = gt expdur(index, /shut dur)

5.4.11 GT EXPLAT

GT EXPLAT will calculate the latency time between the time which is stored in the index and the actual time the integration started. The value returned is in milliseconds and is typically between 50 and 100.

IDL> explat = gt explat(index)

5.4.12 GT EXPMODE

GT EXPMODE routine returns the exposure mode. The values it can return are:

0 Normal shuttered exposure
1 Dark image (shutter stays closed)
2 Calibration image (shutter open during readout)

IDL> expmode = gt expmode(index)

5.4.13 GT CORNER

GT CORNER will return the full resolution IDL (FRE) coordinates of the center of the pixel in the lower left hand corner.

IDL> corners = gt corner(roadmap) IDL> x0 = gt corner(index, /x)

5.4.14 GT CENTER

The output from GT CENTER can be the location of the center of the field of view in (a) full resolution IDL (FRE) pixels, (b) heliocentric coordinates, (c) the angle E/W and N/S from the center of the sun in arcseconds, or (d) the HXT equivalent pixel coordinates. The default is to read the ATT S/C pointing database for the true de-jittered S/C pointing. It is possible to use the commanded S/C pointing information by using the /CMD option, which is much faster and only of slightly reduced accuracy.

IDL> pix = gt center(index) IDL> helio = gt center(index, /helio) IDL> helio = gt center(index, /helio, /cmd) IDL> ang = gt center(roadmap)

where ang(0,*) is arcseconds east/west with east being negative and ang(1,*) is arcseconds north/south with north being positive. helio is in degrees, and pix is in full resolution pixels. It is possible to get the heliocentric coordinates in a string of the form “N12W23” by using the /STRING option.

5.4.15 GT SXT AXIS

The optical axis of the X-ray telescope has been determined from pre-launch measurements of the effective area as a function of off-axis angle. The analysis is described more fully in SXT Calibration Note 36 (R. Fuller, Jan-94). The relative offset between the X-ray optical axis and the optical axis in the Wide band or Narrow band filters was determined by Tom Metcalf (22-Oct-92). The combined results are summarized below:

                      Optical Axis (CCD Full-Res Pixels)
                             x             y
      X-ray                 515           633
      Narrow Band (NaBan)   514.64        634.47
      Wide Band   (WdBan)   515.24        634.26

The values can be quickly acquired on line using GT SXT AXIS. The next three lines will reproduce the values in the table above. Note that the routine also accepts an index or roadmap data structure and that the filter will be determined from the SXT structure.

IDL> print,gt sxt axis(0) ; 0=X-ray IDL> print,gt sxt axis(1) ; 1=NaBan IDL> print,gt sxt axis(2) ; 2=WdBan IDL> print,gt sxt axis(1,/rel) ; NaBan relative to X-ray

The sign convention can be understood if you remember that the NaBan image is Northeast of the X-ray image.

The X-ray optical axis has a somewhat large uncertainty, since the effective area function is broad near the center of the field of view of the telescope. The relative precision between the optical and X-ray axes has been determined to better than 0.4 Full-Res pixels (see the Instrument Guide).

The /rel xray keyword option can be used to return the relative offset to the X-ray axis.

5.4.16 GT PIX SIZE

The angular size of the SXT full-resolution pixel was determined from pre-launch focal length measurements to be 2.45280 arc sec for the X-ray telescope. Half- and quarter-resolution images are twice and four times this value, respectively. A routine to provide this number is GT PIX SIZE which is called as:

IDL> print,gt pix size() IDL> print,gt pix size(index)

If the SXT index or roadmap is passed in, the routine will return the appropriate pixel size in arc sec for the resolution of the data. The pixel size of the optical images is thought to be within 1% of the X-ray pixel size.

5.4.17 GT TEMP CCD

GT TEMP CCD returns the temperature of the CCD in degrees Celsius.

IDL> temp = gt temp ccd(index)

5.4.18 GT TEMP HK

GT TEMP HK returns the temperature measured by the platinum resistance thermometer (PRT) on the spacecraft and SXT instrument. There are 16 temperatures saved in the SXT index.

IDL> temp = gt temp hk(index, 0, title=title)

5.4.19 SXTPNT SUM

The routine SXTPNT SUM will read the observing log and give a summary of the unique PFI pointing for a given time period. It groups the summary listing based on each unique pointing location over time. See the User’s Guide for a sample listing. For the command:

IDL> sxtpnt sum, ’5-oct-93 13:00’, ’6-oct-93 1:00’

it will produce a file in your home directory with the name “sxtpnt sum.txt” (unless you specify the output file name with the outfil keyword option).

5.4.20 PR SXTOBS

The routine PR SXTOBS will read the observing log and list each and every FFI and PFI image taken for the times requested. The format of the output is simply the output of GET INFO which is described in detail in the Reference Guide. See the User’s Guide for a sample listing. The keyword /long changes the formatting and information displayed.

IDL> pr sxtobs,’15-nov-91 22:00’,’15-nov-91 22:05’ IDL> pr sxtobs,’15-nov-91 22:00’,’15-nov-91 22:05’,/long 5.4.21 PFI LOC

PFI LOC will display the SXT PFI location. There are two modes which can be used when calling PFI LOC. If the first parameter is a date and time string, it will read the observing log and make a simple plot showing the location relative to the sun’s limbs. If the first parameter is an SXT index, then it assumes that you have displayed this full frame image, and it reads the observing log and puts a box on the image where the PFI is located. Some sample calls are:

IDL> pfi loc, ’23-jun-93 1:00’ IDL> pfi loc, timstr, hours=2 IDL> pfi loc, index IDL> pfi loc, index, pfi struct=sxtp IDL> pfi loc, index, ’6-nov-93 3:30’ IDL> pfi loc, index, ’6-nov-93 3:30’, /hour

It is possible to use the hours keyword option to specify the number of hours plus and minus the full frame time for which you wish to have PFI locations displayed. The pfi struct option exists so that you can pass a PFI roadmap or index structure directly to the PFI LOC to be displayed over the image. If a second positional parameter is passed, the routine will get the PFI locations for that time and overlay them on the image.

When using STEPPER, the “o” option will call PFI LOC automatically for you, taking into account any rebinning which has been performed on the image.

5.5 Routines for Calibration and Analysis

5.5.1 SXT PREP

SXT PREP can perform a wide variety of the standard processing required to use SXT data. The User’s Guide has a description of the capabilities and some of the fundamentals, and it should be read before reading this section.

The input specifying which images to process can be either an index and data, or it can be the filenames infil and dataset numbers dset arr. Some sample calling sequences are:

IDL> sxt prep, input1, input2, index out, data out IDL> sxt prep, index, data, index out, data out IDL> sxt prep, infil, dset arr, index out, data out, $ IDL> helio=[-9.8,-20.3], date helio=’14-JUN-92 02:37:41’ IDL> sxt prep, index, data, index out, data out, /helio, ref image=10 IDL> sxt prep, index, data, index out, data out, ref image=spr index(0) IDL> sxt prep, index, data, index out, data out, /roll

A description of some of the keywords follows:

——————– Calibration Parameters ——————–

nocorrections If set, then LEAK SUB, and therefore DARK SUB, are not called.

dc scalar If set, then a simple scalar value is used for dark current subtraction.

Default is to take the image closest in exposure.

dc interpolate If set, then perform interpolation on the dark frame images to get a better background subtraction. Default is to take the image closest in exposure.

upper only If set, only flag +1 pixels above (in a column) the saturated area (not

the −1 pixels) [see call to SXT SATPIX].

sfd corr If set, then the input image is an SFD image. Use the SFD decompression, perform the registration, and recompress using the SFD compression.

——————– Alignment Parameters ——————–

register If set, then perform the registration and correct only for the S/C jitter.

Using the HELIO or SUNCOORD options selects this option also.

ref image A single roadmap or index structure which will be used to define the sun or heliocentric coordinates for alignment. It can be the image number (subscript) of the images being extracted which should be used (not the absolute dataset number within the images listed in ”infil”)

helio the reference latitude and longitude for the alignment. This should be the position of the active region at some time, preferably near the time it crossed closest to the center of the disk. helio(0) = longitude E/W (E is negative) helio(1) = latitude N/S (S is negative) It can be a non-zero scalar value if REF IMAGE is passed in. In this case, the heliocentric coordinates are derived from REF IMAGE.

date helio the reference date/time for the heliocentric input.

suncoord The coordinates of the center of the portion to extract in angular arcseconds relative to sun center. suncoord(0) = E–W coordinate (E is negative) suncoord(1) = N–S coordinate (S is negative)

ccdcoord The FRE (full resolution CCD pixel equivalent) of the center of the

portion to extract.

outres The resolution of the output image (0=FR, 1=HR, 2=QR). If not set,

then convert all to full res. Only relevant if registering the images.

outsize The dimension of the output image [#col, #lin]. If it is a scalar value, it will make the #lin = #col. If not set, it will make it the largest size of the input field of views. If PFI and FFI are mixed, it will take the largest PFI. fill gap A gap can exist between two PFI strips for a single observing region

because of a change in S/C pointing. If this variable is set, then the gap between the PFI strips is filled by interpolating the lines above and below.

roll do  If set, then perform the corrections for the roll offset.
normalize If set, perform exposure normalization.

——————– Miscellaneous Parameters ———-

force read If set, read all datasets at once even if input was specified with IN-FIL/DSET. This is available because doing LEAK SUB on all images at once if the input is PFI is more efficient.

outfil If specified, the data will not be passed back in INDEX OUT and DATA OUT. Instead, it will be saved in an SDA file with the name specified by this parameter. If “uncert” and “satpix” are included in the procedure call, they will be saved with the same file name with a “ unc” and a “ sat” appended to the file name.

SXT PREP always stores the actual pixel address of the portion extracted and used. The .HIS.SUN CENTER value stored in the index is always the actual pixel address where the sun center fell on the CCD for that time. When a registration is performed, the CORNER SAV value will not be set to the same value, even though it is a registered cube. This is because the registration is RELATIVE TO SUN CENTER, and that is changing. Plotting the corner minus the sun center WILL result in a flat plot with all values identical.

5.5.2 XSXT PREP

XSXT PREP was under development as this volume was being written. See the User’s Guide for a overview of the capabilities of XSXT PREP.

5.5.3 SXT DECOMP

The SXT data are normally compressed from a 12-bit word to an 8-bit word via a pseudosquare-root compression algorithm. SXT DECOMP takes the compressed data array (data) and creates an integer array (ndat) of the decompressed numbers by typing:

IDL> ndat = sxt decomp(data)

5.5.4 SXT COMP

It is possible to re-compress the data. The routine SXT COMP will compress integer data which is in the range between 0 and 4095 into byte data using the SXT compression algorithm.

IDL> bdat = sxt comp(data)

5.5.5 RESTORE LOW8

RESTORE LOW8 will combine a compressed image with a low 8-bit image to restore the full 12 bits of resolution. The images need to be taken relatively close in time (less than hours apart), and it must be for the same PFI location. A sample call is

IDL> out = restore low8(low8, comp)

where low8 is the SXT image taken with the low 8 bits, and comp is the comparable compressed image (ideally, acquired at nearly the same time).

5.5.6 SFD COMP

The SFD (SXT Full-frame Desaturated) images use a logarithmic compression in order to save the data as byte type. SXT COMP compresses the data into byte type. The compression algorithm assumes that the data input is in DN/sec/HR pixel and the values range between 0 and 10^{6}.

IDL> data out = sfd comp(data in)

5.5.7 SFD DECOMP

The SFD (SXT Full-frame Desaturated) images use a logarithmic compression in order to save the data as byte type. The decompression algorithm is data out = 10.^{(data in/255.∗6)}. The routine SFD DECOMP will do this decompression for you. Beware: the output are floating type numbers and therefore, require four times as much memory.

IDL> data out = sfd decomp(data in)

5.5.8 EXP NORM

EXP NORM ^{2} takes a data array and produces an exposure and summation mode normalized floating point array. Note the floating array takes four times more room so don’t use too big an array to start with. You need to give it an index array – make sure that the two correspond exactly or else the results will be meaningless. To run it, type:

IDL> ndat = exp norm(data,index) IDL> ndat = exp norm(data,index,0)

2If this code is used as a template, we recommend that the calling sequence be index,data for all routines

To do this it follows the following steps:

• decompresses the data array

• subtracts background

• divides by the exposure time

• divides by the number of pixels binned

5.5.9 SXT COMPOSITE

SXT COMPOSITE combines two or three SXT images with different exposures into a “composite” image. The longest exposure is used as the basis for the final exposure. SXT COMPOSITE uses the results of SXT SATPIX to identify the locations of saturated pixels and replaces them with equivalent pixels from the shorter exposure images, scaled for the long exposure image. This routine uses the algorithm that is used to produce the “sfd” images in the standard SXT movie. Example calls are:

IDL> cimg = sxt composite(index,data) IDL> cimg = sxt composite(index,data,satpix=satpix) IDL> cimg = sxt composite(index,data,dc data,sx=[1,2,3],sd=[0,2,1])

By default, SXT COMPOSITE will do dark and leak frame subtraction if data is byte-type. The user may optionally specify the dark-frame images in dc data and will default to data − dc data unless the explicit X-ray and Dark-frame indices are given in sx and sd, respectively.

Note that the replaced pixels are box-car smoothed with a smoothing window of 5 pixels.

5.5.10 MK SFD

MK SFD creates a desaturated set of images from a series of long and short exposures by typing:

IDL> mk sfd,infil,outfil,filpref

where infil is a string array of one or more SFR filenames (and locations), outfil is the string array containing the name of directory where you want the file to be created, and filpref is a string array which contains the three letter file prefix name (usually SFD). The desaturated image is compressed such that:

                      0 =         1 dn/sec/5 arcsec pixel
                    255 = 1,000,000 dn/sec/5 arcsec pixel

Use the routine SFD DECOMP to restore the data from compressed bytes to real numbers.

5.5.11 DARK SUB

It is possible to have the dark current removed from PFI and FFI images by using the following command:

IDL> data out = dark sub(index,data)

The routine can interpolate between two reference images to simulate the input exposure duration by using the /interpolate option. It will call the routine GET DC IMAGE to get the proper image and take the subsection for the PFI cases. The SDC data files for the time period being calibrated must be in the directory $DIR SXT SDC.

The /interpolate option can be turned on and run automatically every time DARK SUB is run by setting the environment variable ys dark interp. An example of how to turn this on and off is:

% setenv ys dark interp 1 # will enable automatic interplation % unsetenv ys dark interp # will disable automatic interplation

LEAK SUB calls DARK SUB if the input data is BYTE type, so normally a user does not need to call DARK SUB directly.

5.5.12 GET DC IMAGE

These routines get an appropriate dark frame (dcdata) and its index (dcindex) to match the images you are working with. The routines find the full-frame dark image that is closest in time and exposure level. The routine can interpolate between two reference images to simulate the input exposure duration by using the /INTERPOLATE option. It can be called with the index for a single image by typing:

IDL> get dc image, index(3), dcindex, dcdata

where index(3) is the index of the image you are working with, or you can pass a list array of indicies by typing:

IDL> get dc image, index, dcindex, dcdata, imap

For this case, imap is the same length as index and tells you which image to use in dcdata for the corresponding index. For example, if index is 10 elements long, and imap(8)=3, then you should use dcdata(*,*,3) for the image that goes with index(8).

It is possible to extract the dark frames manually using a command similar to the following example:

IDL> get dc image, xxx,dc index, dc data, times=’1-sep-92 1:00’, res=0, dpe=15 The SDC data files for the time period being calibrated must be in the directory $DIR SXT SDC (and the SDW data files must be in $DIR SXT SDW when calibrating images which were taken when the TEC is off).

5.5.13 LEAK SUB

On 13-Nov-92, there was an entrance filter failure which caused visible light to contaminate the X-ray images. It is possible to have the leak image removed from PFI and FFI images by using the following command:

IDL> data out = leak sub(index,data)

The routine will call DARK SUB if it comes in as byte type. The technique at the time of writing this guide was to find the sunset leak image which matched the following criteria

• The same X-ray analysis filter.

• Using the ND filter if the input X-ray image used the ND filter.

• The nominal S/C pointing must be the same (there is one set of images between 13-Nov-92 and 19-Aug-93, and another set for images taken after 19-Aug-93)

The small changes in S/C pointing were not being accounted for as of Jan-94, even though there is a relatively strong dependence on pointing drift. Future modifications to LEAK SUB might take this factor into account.

The leak image files must be in the directory $DIR SXT SFC.

5.5.14 INTERP IMG

To linearly interpolate two images to a specified time type:

IDL> interp img, index1, img1, index2, data2, time, index out, img out IDL> interp img, idx(10),img(*,*,10),idx(12),img(*,*,12), time, nidx, nimg

where the two input images to be interpolated are specified by the paired variables index and img. The input interpolation variable time can be in any Yohkoh standard time format. The returned parameters index out and img out refer to the interpolated index information and the interpolated image data respectively.

5.5.15 SXT FSP and SXTBOX FSP [*]

SXT FSP is the same as HXS FSP, but for SXT spectra, you now have the option to fit temperatures and emission measures using more than two filters.

IDL> SXT FSP, index, data, fit pars SXTBOX FSP fits spectra to a given box, chosen by LCUR IMAGE, for a set of SXT images.

IDL> SXTBOX FSP, index, data, fit pars [, sc par, ch dta, sdel=sdel, $ IDL> outfile=outfile, pfile=pfile, same bx=same bx, boxq=boxq]

5.5.16 SXTHXT FSP and SXTHXTBOX FSP [*]

See the description under in the HXT chapter.

5.5.17 SXT TEEM

SXT TEEM enables you to derive a temperature map (Te) from two filters taken of the same region. Some sample calls are:

IDL> sxt teem,index1,data1,index2,data2,Te IDL> sxt teem,index1,data1,index2,data2,Te,EM,dTe,dEM,subs=subs IDL> sxt teem,index1,data1,index2,data2,Te,EM,dTe,dEM,$ IDL> sat1=sat1,sat2=sat2,unc1=unc1,unc2=unc2 IDL> sxt teem,index1,data1,index2,data2,Te,EM,dTe,dEM,$ IDL> sat1=sat1,sat2=sat2,unc1=unc1,unc2=unc2,/average

index1,data1 should contain data for a single SXT filter. index2,data2 should contain data for a single SXT filter which is different from that contained in index1 and data1.

The input data should be decompressed and background subtracted. If index is a vector, then the data arrays should be 3-d arrays (cubes).

Normally, SXT TEEM will get the proper exposure times from the index. However, if you have renormalized the data in a non-standard manner, then use the t1=t1 and t2=t2 keyword to over-ride the value in index. t1 = Integration time in msec. If Image1 is a cube, then t1 must be a vector and whose length = n elements(Image1(0,0,*))

The keyword option sum = sum calls SXT SUMXY to rebin the data. The input image must be a 2-d or 3-d array sum must be an integral fraction of the number of X and Y elements. If either of the two preceding conditions is not met, summing is not done.

If data1 is a cube, Te will be computed for each image by default. If subs=subs is present, subs should contain the indices of image pairs to analyze.

The keyword option thresh1=thresh1, thresh2=thresh2 sets the level above which data are considered for processing. The level is set on the background subtracted, rebinned data values. If sum=2, then you must increase the value of thresh1 by a factor of 4 to have the same effect.

5.5.18 SXT ERG PER DN

SXT ERG PER DN returns the log of the radiative energy per DN assuming a coronal spectral model with temperature, T_{e}. The units of the result are log_{10}(erg/DN), where DN is assumed to be the decompressed and background subtracted value that is read out of the CCD camera. Example calling statements are:

IDL> log10 erg dn = sxt erg per dn(log10 Te,filter) IDL> log10 erg dn = sxt erg per dn(log10 Te,filter,date=’1-dec-92’)

and log10 T e and filter are input variables which may be vectors. log10 T e is the log of the electron temperature (K), of the solar plasma and filter is the SXT filter number. For example, to compute log 10(erg/DN) for all filters over the range of temperatures between log 10(T_{e}) between 5.5 and 8.0 use the following calling sequence:

IDL> Te = 5.5 + findgen(26)*.1 ; Generate a temperature vector IDL> log10 erg dn = sxt erg per dn(Te,indgen(6)+1)

The purpose of the date keyword option is to allow the post-November 1992 SXT response functions to be used. This can be invoked by specifying any date later than “14-Nov-92”, the day after a hole in the entrance filter appeared, and caused the SXT response function to change.

5.5.19 SXT VIGNETTE

The SXT vignette function was determined from pre-launch calibration data obtained at White Sands Missile Range (WSMR) and was further refined using flight data. The vignette function is approximated by two non-concentric cones. The routine SXT VIGNETTE computes the SXT vignette function for a given SXT image as described in the scalar SXT index as follows:

IDL> vignette = sxt vignette(index) ; SXT Index should be a scalar IDL> vignette = sxt vignette(array) IDL> vignette = sxt vignette(index,energy)

If array is given then array = [x0, y0, Nx, Ny, Mode] and Mode = 0, 1, or 2, for full, half and quarter resolution. The second argument, energy will specify whether to use the low-energy results (derived from C-K and Al-K test data) or the high-energy results (derived from Ag-L test data). SXT VIGNETTE is called by SXT OFF AXIS to correct SXT images. A quick summary is available by typing:

IDL> help vignette ; Display SXT vignette function 5.5.20 SXT OFF AXIS

The X-ray vignetting function can be approximately removed with the routine SXT OFF AXIS. Example calling statements are:

IDL> data out = sxt off axis(index, data, index out) IDL> data out = sxt off axis(index, data, /update index) IDL> vignette = sxt off axis(index, /vignette only)

SXT OFF AXIS calls SXT VIGNETTE to calculate the SXT X-ray vignette function. See SXT Calibration Note 37 and the Instrument Guide of this manual for more information. Two Vignette functions have been determined: the high-energy function is applied to the Al12 and Be119 filters and the low-energy function is applied to all other filters. Note that this approach is somewhat approximate and care should be used when applying this correction to SXT data.

5.5.21 WL PROCESS [*]

Alan McAllister wrote the routine WL PROCESS to process the SXT aspect sensor images and generate limb darkening profiles, flat field images, white light images minus the dark current, flat field, and limb darkening signals. The routine is still under development and you should look at the documentation header of that routine for more details.

5.6 Routines for Image Alignment

The routines described here are specific to SXT. There are also some routines listed in the General Yohkoh Software chapter which can be used for either SXT or other ground-based images.

5.6.1 SXT PREP

See the earlier section describing the details of SXT PREP.

5.6.2 SXT CENTER [*]

SXT CENTER calculates the row (x) and column (y) of the center of the solar disk and the radius (r) of the Sun given a data cube (data) and the index structure (index). The routine performs a fit to the limbs of the image. The fit can be off by several pixels in the E/W direction for X-ray images. See the SXC database if you are interested in the X-ray center for all SFD images as calculated by SXT CENTER.

IDL> sxt center,data,index,x,y,r

There are many options that are possible to use, please check them with DOC LIBRARY.

5.6.3 GET GBO PFI [*]

Given a SXT PFI image and either the file name of a ground-based image, or the ground-based image itself, GET GBO PFI will extract the portion of the GBO image corresponding to the SXT PFI. The routine (a modification of the routine GBO PFI written by Alan McAllister) determines the PFI corners from index, rotates these to the time of the GBO image, and extracts the section of the GBO image. You can either pass in the GBO image itself, in which case you MUST also supply the time for it, or a string with the path/filename using the Yohkoh file name convention. Some sample calls are:

IDL> gbo img = get gbo pfi(index, gbo fullimg, gbo time) IDL> gbo img = get gbo pfi(index, ’/yd5/gbo91 46a/gkm911116.1958’)

5.6.4 OCENTER

OCENTER will obtain the center of the Sun from a optical full frame image. It makes a “silhouette image” consisting of 1’s and 0’s, then calculates the centroid to determine x and y. The area is used to determine the image radius. The input image can be compressed or decompressed, and in any summation mode. The output is in SXT full resolution pixels unless /noscale is set.

IDL> ocenter, image, x, y, radius IDL> ocenter, image, x, y, radius, /noscale, cut=0.3

where image is the 2-D input image, and x and y are the E/W and N/S position of the sun. radius is the solar radius in pixels and cut is the fraction of image maximum to define the limb (default is 0.2)

5.7 Routines for Accessing Secondary Database Files

5.7.1 GTAB COMM, GTAB ENTRY, GTAB ROI

It is possible to get information on what was used in an SXT table by using one of these routines. GTAB COMM returns information on the common table, GTAB ENTRY returns information on the entry tables, and GTAB ROI returns information on the ROI location and image shape tables. Some sample calls:

IDL> print, gtab comm(index)

IDL> print, gtab comm(’15-nov-91’)

5.7.2 GTAB PFI and GTAB FFI

These routines allow a user to print information on the partial-frame and full-frame observing sequence being used. There are four possible sequences for PFI and four more for FFI. Which sequence is used is determined by the DP mode and telemetry rate. If the SXT index for an image is passed, then it will figure out which of the four sequences was running at that time.

IDL> print, gtab ffi(index) IDL> print, gtab pfi(’15-nov-91’, 2)

If a time is passed, it will default to show sequence number 0. It is possible to specify which sequence table by passing it as a second parameter.

5.8 Routines for Engineering Plots

5.8.1 PLOT SOT

PLOT SOT will plot out the aspect sensor degradation as a function of time by typing:

IDL> .run plot sot

It will produce plots for both the narrow-band and wide-band filters.

5.8.2 PLOT TEMPS2 [*]

PLOT TEMPS2 plots the SXT instrument temperatures. This program can only be run after the variable index exists which came from an SXT file. A common practice is to run PLOT SOT which gets images over the whole mission.

IDL> .run plot temps2

A mosaic of time histories of the various instrument temperatures will be plotted.

5.8.3 PLOT SSL

PLOT SSL will plot the SXT summary log (SSL) information. A series of plots are made which include:

• Plots of measured exposure levels

• Plots of SXT temperatures

• Plots of number and duration of exposures through different filters

• Plots of percentage of data received 5.8 Routines for Engineering Plots 63

• Plots of average/max SXS1,SXS2,HXT-Low,RBM counting rates Sample calls are:

IDL> plot ssl IDL> plot ssl, [0,1] IDL> plot ssl, 3 where the second example only plots the first two series of plots (exposure levels and SXT temperatures), and the last example only plots percentage of data received.

6 Wide Band Spectrometer (WBS)

6.1 Routines for Making Time Plots

6.1.1 PLOTT WDA PLOTT WDA will make a light curve plot for all channels using either the index or roadmap. The second ‘T’ signifies that it is a time plotting routine. Seven plots are made, one for each channel. Some sample calls are:

IDL> plott wda, index IDL> plott wda, roadmap IDL> plott wda, roadmap(100:200), psym=10

6.2 Routines for Making Spectral Plots

6.2.1 PLOTS WDA PLOTS WDA is a very rudimentary plotting routine that will display the spectra for any of the WBS sub-instruments. It is interactive and will prompt the user for which sub-instrument to display, and which datasets. A sample call is:

IDL> plots wda, index, data

6.3 Routines for 2-D Spectral Display

6.3.1 DISP WDA The evolution of all of the WBS spectra against time can be displayed as a pseudo-image using the routine DISP WDA. Note: The time axis of DISP WDA is not uniform.

IDL> disp wda, index, data

6.3.2 SXSPC SXSPC makes the time profiles of SXS-PC11, PC12, PC-21, and PC22 with various options^{3}.

IDL> pc data= sxscv(data, index) IDL> sxspc,pc data,index 6.3.3 WBSPC WBSPC makes four counting rate time profiles of SXS-PC11(21), SXS-PC12(22), HXS-PC1 and HXS-PC2.

3If this code is used as a template, we recommend that the calling sequence be index,data for all routines

IDL> wbspc,data,index

6.3.4 PLOT HXSPC PLOT HXSPC makes a counting rate time profile of HXS-PC1, HXS-PC2 or HXS-PC1 + PC2 with UTPLOT.

IDL> plot hxspc,index,data,pc

6.3.5 PLOT HXSPH PLOT HXSPH makes a counting rate time profile of HXS-PH with UTPLOT. You choose the energy or channel range.

IDL> plot hxsph,index,data,range=range IDL> plot hxsph,index,data,channel=channel

6.3.6 PLOT GRSPCL PLOT GRSPCL makes a counting rate time profile of GRS-PC11, PC12, PC21 or PC22 with UTPLOT.

IDL> plot grspcl,index,data,pc

6.3.7 PLOT GRSPCH PLOT GRSPCH makes a counting rate time profile of GRS-PC13, PC14, PC15, PC16, PC23, PC24, PC25 or PC26 with UTPLOT.

IDL> plot grspch,index,data,pc

6.3.8 PLOT GRSPHL PLOT GRSPHL makes a counting rate time profile of GRS-PHL1 or GRS-PHL2 with UTPLOT. You choose the energy or channel range.

IDL> plot grsphl,index,data,ph=1,range=range,channel=channel 6.3.9 PLOT GRSPHH PLOT GRSPHH makes a counting rate time profile of GRS-PHH1 or GRS-PHH2 with UTPLOT. You choose the energy or channel range.

IDL> plot grsphh,index,data,ph=1,range=range,channel=channel

6.3.10 Light Curve of RBM-PC See the description in the User’s Guide.

6.4 Routines for Data and Information Extraction

6.4.1 GT Routines The routines GT SXS1, GT SXS2, GT HXS, GT GRS1, GT GRS2, GT RBMSD, and GT RBMSC routines will extract the average counts/sec for the selected channel.

SXS1  Only SXS PC21  3-15 keV
SXS2  Only SXS PC22  15-40 keV
HXS  HXS PC1 plus HXS PC2  20-657 keV
GRS1  GRS PC11 plus GRS PC21 0.3-1.2 MeV
GRS2  GRS PC12 plus GRS PC22 1.2-5.6 MeV
RBMSD PC1 plus PC2  5-300 keV
RBMSC Only RMSSC  >20 keV

The input can be roadmap, index, or observing log and it will get the proper structure tag and decompress it properly to return counts/sec/sensor. It is possible to get a string defining the channel selected by using the title keyword option.

IDL> y = gt sxs1(roadmap) IDL> y = gt hxs(w h, title=title)

6.4.2 GET INFO See the description in the “General Yohkoh Software” chapter on page 17.

6.5 Routines for Calibration and Analysis

6.5.1 HXS FSP [*] With HXS FSP you interactively choose data intervals, and fit the HXS spectrum, using the routine fsp 11.pro. index and data are the WBS index and data structures. fit pars is the structure containing the results. Optional parameters are – sc par is a structure containing spectrometer channel information, ch dta is a structure containing the data for each channel, outfile and pfile are filenames for printouts and plots, respectively; countfile is a filename, if this option is set, the count rates are printed into that file which may be fit by the routine FSP PROC described above.

IDL> HXS FSP, index, data, fit pars IDL> HXS FSP, index, data, fit pars, sc par, ch dta IDL> HXS FSP, index, data, outfile=outfile IDL> HXS FSP, index, data, pfile=pfile IDL> HXS FSP, index, data, countfile=countfile IDL> HXS FSP, index, data, fit pars, sc par, ch dta, sdel=sdel

6.5.2 HXS SPL

HXS SPL makes a HXS spectrum fitted by single power law. You interactively choose the background and flare time intervals for the spectral analysis.

IDL> hxs spl,data,index

6.5.3 HXS SP2

HXS SP2 makes HXS spectrum fitted by double power law in a similar way to HXS SP1.

IDL> hxs sp2,data,index

6.5.4 GRS SPEFF

GRS SPEFF makes preliminary GRS-1 and 2 spectra (0.3 – 100 MeV).You interactively choose the flare and background time intervals for the spectral analysis. The flux of each channel is calculated by dividing the counts by full energy absorption efficiency. The first three channels data of GRS-1 and the four channels of data of GRS-2 are not plotted because these channel widths are not exactly determined.

IDL> grs speff,data,index,ph=1,/disp IDL> grs speff,data,index,ph=2,/disp IDL> grs speff,data,index,ph=1,fltime=fltime,bgtime=bgtime

6.5.5 Further Analysis of HXS and GRS Data

There are several routines for further HXS and GRS data analysis. With these routines you can get the following:

• Time profile of ratio of HXS-PC2 to HXS-PC1 counting rates

• Time profile of ratio of HXS-PHi-j to HXS-PHm-n counting rates

• Time profile of ratio of GRS-PCi to GRS-PCj counting rates

• Time profile of ratio of GRS-PHi-j to GRS-PHm-n counting rates

• HXS spectral fitting by single power law, double power law, thermal and power law plus Gaussian etc.

• Residuals and chi-square map for HXS spectral analysis

• Gamma-ray spectral analysis (continuum plus lines) with the maximum number of fitting parameters is 62

• Residuals and chi-square map for GRS spectral analysis

See the detailed description in the User’s Guide.

6.6 Instrument Response Files

The instrument response files for SXS, HXS and GRS are listed here. These files are ASCII files and are located in $DIR WBS CAL.

6.6.1 SXS

dtcorrf.dat Table of the deadtime correction factor vs. counting rates of SXS-PC11, PC12, PC21 and PC22 6.6.2 HXS

hxs 01.rel Table of channel number (0 – 31) vs. energy (keV) relation for HXS- PH for HXS-OS of “01 (hex)”. (1992 June 9 – present).

hxs 21.rel Table of channel number vs. energy relation for HXS-PH

for HXS-OS of “21 (hex)”. (1991 Oct. 1 – 1992 June 9).

hxs 01 conv260.rel Table of incident photon energies (keV) corresponding to the 260 rows of the 260 x 32 response matrix for HXS-PH spectral analysis for HXS-OS of “01 (hex)”. (1992 June 9 – present).

hxs 21 conv260.rel Table of incident photon energies (keV) corresponding to the 260 rows of the 260 x 32 response matrix for HXS-PH spectral analysis for HXS-OS of “21 (hex)”. (1991 Oct. 1 – 1992 June 9).

hxs 01 conv260.resp 260 (row) x 32 (column) response matrix for HXS-PH spectral analysis for HXS-OS of “01 (hex)”. The row and column represent the incident photon energy and observed channel number, respectively. The matrix elements of i-th row represent a response function for incident photons of i-th energy in hxs conv260.rel.

hxs 21 conv260.resp 260 (row) x 32 (column) response matrix for HXS-PH spectral analysis for HXS-OS of “21 (hex)”. This is similar to hxs 01 conv260.resp.

dtcf pc hxs.pro Deadtime correction factor for HXS-PC1 and PC2. The correction is not exact when the deadtime-corrected count of total PC exceeds about 100,000 counts/sec. It is due to instrument limitation.

dtcf ph hxs.pro Deadtime correction factor for HXS-PH. The correction is not exact when the deadtime-corrected count of total PH exceeds about 35,000 counts/sec. In addition, spectral distorsion occurs when the deadtime-corrected count of total PH exceeds about 21,000 counts/sec. It is due to instrument limitation.

6.6.3 GRS

grs1 40.rel Table of channel number (0 – 127) vs. energy (keV) relation for GRS- 1 for GRS1-OS of “40 (hex)”. (1991 Oct.1present). Energy widths of the first 3 channels are not exact.

grs2 40.rel Table of channel number (0 – 127) vs. energy (keV) relation for GRS-2 for GRS2-OS of “40 (hex)”. (1991 Oct.1present). Energy widths of first 4 channels are not exact.

grs1 40 conv260.rel Table of incident photon energies (keV) corresponding to the 260 rows of the 260 (row) x 128 (column) response matrix for GRS- PHL1 spectral analysis.

grs1 40 conv260.resp 260 (row) x 128 (column) response matrix for GRS-PHL1 spectral analysis. The row and column represent the incident photon energies and observed channel number, respectively. The matrix elements of i-th row represent a response function for incident photon of i-th energy in grs1 conv260.rel.

grs2 40 conv260.rel Table of incident photon energies (keV) corresponding to the 260 rows of the 260 (row) x 128 (column) response matrix for GRS- PHL2 spectral analysis.

grs2 40 conv260.resp 260 (row) x 128 (column) response matrix for GRS-PHL2

spectral analysis. This is similar to grs1 conv260.resp.

dtcf ph grs11.pro Deadtime correction factor for GRS-PHL1. dtcf ph grs21.pro Deadtime correction factor for GRS-PHL2.

A Yohkoh Distribution and Archive Tapes

A.1 The Organization of the Tapes

A.2 Overview of Yohkoh Data Archive

The YOHKOH data archive is created on Unix computers at ISAS, Japan, using standard archiving tools from the Unix operating system. All archiving is done on 8mm Exabyte tapes. Nominally, each archive tape will contain one calendar week (always beginning with Sunday) of reformatted data and all the software used to access the data. Tapes are created as soon as a complete week of the data are assembled; but, because data are collected from two sources (KSC in Japan and NASA’s DSN) some delay is expected. The tapes are then sent to California for duplication and distribution to the U.S. Co-I’s. At ISAS the archive tapes are copied to MO disk.

Archive tapes can be read on Unix and VMS machines such as SUN, SGI, and Digital using software package written in Interactive Data Language (IDL).

For more information on reading archive tapes see the GO RDTAP description in the User’s Guide.

A.3 Structure of Archive Tape

YOHKOH archive tapes contain a tape directory, reformatted data and software. Tapes are created on DIGITAL’s Unix (Ultrix) system using ‘cp’, ‘cpio’ and ‘tar’ archiving utilities. Reformatted data files and the weekly files are all copied to tape using the ‘cpio’ utility. The software archive is copied to tape with the ‘tar’ utility. Since, the software archive comprises several different directory trees the ‘tar’ utility was the logical choice because it allows one preserve these directory structures during restore operations. Archive tapes will nominally contain a single calendar week (always beginning with Sunday) of reformatted data. (Where each reformatted file contains data from a single orbit which spans approximately 90 minutes.) This calendar week is given a number (week number: 1–53) which is then encoded into the name of the tape in the format ‘yy wka.xx’, where ‘yy’ is the year, ‘wk’ is the week number, ‘a’ is the media number (always “a” for tapes, but can be “b” for other media, such as MO disks), ‘xx’ is the revision number or level of the archive tape.

The detailed contents of a YOHKOH data archive tape begins with two directory files in different formats (ASCII and binary), followed by a series of ‘cpio’ data archive files, and lastly a ‘tar’ software archive file.

YOHKOH ARCHIVE TAPE

TAR Software Archive Files

[figure] [figure]

Archive file 3 with reformatted files: ada, cba, bda, hda, and wda.

Archive file 2 with reformatted files: ada, cba, bda, hda, sfr, spr, and wda.

Archive file 1 with weekly files

[figure] [figure]

CPIO Archive Files (ASCII format for cpio headers)

Tape Directory File Binary format Tape Directory File ASCII format

Structure of the data and software archive tape. Each cpio archive file contains only those reformatted data that share the same file ID (e.g. archive file 2 contains files from the file ID:

911201.1545, file 3 would contain files with file ID 911201.1715).

The tape directory file lists the contents of each tape. The file contains general information such as when the tape was made, the name of the tape, the number of ‘cpio’ archive files, the first and last data file identification (fileID) contained within the ‘cpio’ archive, and a breakdown on the contents of the software ‘tar’ archive file giving names and sizes of embedded ‘tar’ files. The tape directory also has two detailed tables. The first of which tabulates the weekly files present within the first ‘cpio’ archive and the second tabulates the ‘orbit’ fileID and the sizes of each reformatted file (e.g. ada, bda, cba, hda, sfr, spr, wda) for each ‘cpio’ archive file. This last table is very useful for identifying whether a given file is on a tape.

The reformatted data are placed on tape with the help of the Unix archiving utility ‘cpio’ which allows a specified list of files to inserted into a single archive file. Using this utility we have bundled the data by placing files with the same data file identification or fileID (such as 911201.1654, which corresponds to the date 1-Dec-91 at 16 hours and 54 minutes) together into a single archive file and copied to tape. Each fileID is unique and represents one orbit of data. Since there are about 16 orbits per day, each tape could have 112 + 1 different ‘cpio’ archive files. The additional archive file is for the first ‘cpio’ file which only contains the weekly files such as the observing log file (obs*) and the pointing file (pnt*).

Even though the software is placed on the tape, it is best to obtain the software across the network. See Appendix ?? for more details.

A.4 Access Software

There are three IDL utility programs which access the archive tapes. An IDL procedure called MKTAP to create archive tapes, another procedure called RDTAP to read any of the various sections or parts of sections of the tape (see Figure), and finally a robust procedure called YOHKOH SW to read and restore just the software archive files. The YOHKOH SW procedure is very useful and is recommended for creating a software environment which would be quite similar to the ISAS computing environment. Documentation for RDTAP and MKTAP are given below and the ’yohkoh sw’ documentation can be found in separate guide which is attached to this users guide.

A.4.1 Reading Tapes (RDTAP)

When using RDTAP to read and extract data from the archive tapes be aware that the newly created disk files are always placed in the users default directory. RDTAP has an on-line brief help information listing to aid the unfamiliar user. To activate this feature type RDTAP without any parameters at the IDL prompt. To get a listing of the contents of the tape one can either use the RDTAP utility with the option to extract the tape directory files

IDL> rdtap, device, /dir

which creates the two files called ‘xadyy wkd.xx’ (ASCII) and ‘xbdyy wkd.xx’ (binary). An alternative method is to simply copy the first file (the ASCII file) from the tape to disk via DCL or Unix commands (e.g. for VMS mount/for mt: and copy mt: tapedir.lis or for Unix cp dev tapedir.lis). When extracting the binary data files from tape to disk the files are restored without any byte swapping. This keeps the files in their original format so that any byte swapping, if necessary, is handled automatically and transparently to the user at the time the disk files are read by the analysis software. Note that the tape is rewound at the beginning of each RDTAP request, so it is more efficient to specify all requests within a single RDTAP command than it is to make several smaller requests. Also the requested list of the files need not be in order of appearance on the tape as the routine will sort all input requests for the most efficient extraction.

The general calling sequence for RDTAP is:

IDL> rdtap, dev [, file, prefix= prefix, fileID= fileID, week=week, $ IDL> dir=dir, range= range, software= software]

Where the optional parameters are enclosed within brackets (note that when calling RDTAP with the dev parameter specified at least one other optional parameter must also be included. In general, all non-switch parameter inputs have a string data-type (e.g. dev = ‘/dev/nrmt0h’). The definitions of the parameters is as follows:

        dev     is the device name (Ultrix example: '/dev/nrmt0h').
        file    is a list of requested file names (e.g. spr910925.1020).
        prefix  is a list of requested file prefixes.
        fileID  is a list of requested fileIDs (e.g. 910925.1020).
        week    is a list of weekly files to restore.
        dir     is a switch to restore the tape directory file.
        range   is a switch to create a list of files from input.
        soft*   is a switch to restore the software tar file to disk.

Some examples are:

IDL> rdtap, dev, [’spr911029.0300’,’bda911030.0440’] IDL> rdtap, dev, prefix=[’spr’,’sfr’] or rdtap, dev, prefix=[’s’] IDL> rdtap, dev, fileID=[’911029.0900’] IDL> rdtap, dev, prefix=’spr’,fileID=’910923.1234’, week=’all’ IDL> rdtap, dev, prefix=’s’, fileid=[’911101.0200’,’9111015.0900’],/range IDL> rdtap, def tapd(1), /software, /dir

The first example restores just the requested spr and bda files. The second example restores ALL SXT files from the tape (could take about two hours). The third example restores all archived files of this fileID. The fourth example restores just the spr910923.1234 file and all weekly files. Note, that week=[‘obs’, ‘pnt’, ‘fem’, ‘evn’] will restore any of these files found on the tape (i.e. if FEM and EVN files were missing, only OBS and PNT files would be restored. The fifth example restores sfr and spr files between orbits dated 1-Nov at 02:00 UT and 15-Nov at 09:00 UT. The last example decompresses the software archive file and restores the tape directory files to the default directory. Using the DEF TAPD function to specify the device name by selecting the device number is a simple way to get the correct device name.

In using the tape directory file to review tape contents and select data files to be extracted from the tape, find the final table within the file which tabulates the fileID and file prefix. The heading of table reads FileID and cpio, followed by all of the valid file prefixes. To find a given file with the table, match the fileID with the corresponding row within the FileID column and then match the file prefix with the corresponding prefix column. A non-zero entry for the size of the file (bytes) indicates that the file is on the tape.

A.4.2 Tape Reading on VMS

For VMS systems, there are two DCL commands which must be done prior to entering IDL in order to access the tape drive. Issue the following two commands before starting IDL:

       $ Allocate tapeDrive mt*        where tapeDrive refers to the 'mu'
                                        device (example: 'mub0') and mt*
                                        refers to mt0 thru mt9, pick one.
       $ Mount/FOREIGN mt*     this mounts the tape to the 'logical device'
                               name mt* (again mt0 thru mt9).

Then enter IDL. Assuming one selected ‘mt0’ as the logical device name for the tape drive in the allocation command, the valid rdtap command restoring the tape directory files would look like:

IDL> rdtap, def tapd(0), /dir

A.4.3 Tape Read Errors

Any tape read errors encountered while reading the tape are flagged to the current session and an attempt is made to continue reading the tape. All file(s) which were not restored as a result of the tape read errors are logged to a tape error log file called ‘rdtape.log’.

In general, recovering unread file(s) may be difficult. If an error resulted in the loss of two or more consecutive files one can first attempt to restart the file extraction beginning with the second file in the ‘rdtape.log’. If this procedure is successful, it is most likely that the first file will be unreadable, however one should still attempt to re-read the file as the error may have been due to dust or other contaminates on the tape. If the error log contains several of these situations of two or more consecutive file losses repeat the above procedure for each case.

A.4.4 Writing Tapes

Tapes are created with the IDL procedure MKTAP. A blank 8mm tape should be in the desired tape drive before one starts the procedure. The nominal mode starting this procedure is to specify the device as a parameter:

IDL> mktap, dev

where, dev can be either a string-type variable such as ‘/dev/nrmt0h’ or an integer 0, 1, 2, etc., which specifies the device number. The procedure can also be started with out a device specified. However, in this case the program assume the tape is in one of the follow defaults devices (system dependent): ‘/dev/nrmt0h’ for Ultrix systems, ‘/dev/rmt/tps0d6nrnsv’ for IRIX systems, and ‘/dev/nrst0’ for SUN systems. The first question one must answer is the name of the data directory (note this is NOT the path). For example, if the path to the archive data is ‘/yd4/92 12a’ then the data directory is ‘92 12a’. The second and typically last question asks whether to archive all files or part of the files within the data directory. After answering this final question a script file called ‘mktape.com’ and two tape directory files xbdyy wka.xx (binary version) and xadyy wka.xx (ascii version) are created in the data directory. It is important to note that the file ID of the tape directory files (xad and xbd) is the same as the name of the tape which will hold these files. This completes the IDL procedure MKTAP. To make the tape one types the following command:

% source mktap.com

MKTAP also has a number of switches which can be specified with the MKTAP command to allow the program to run in different modes. Listed below are a few examples, see the program header to MKTAP for a complete list:

IDL> mktap, 1, datadir=’91 50a,91 51a’

will use tape drive ‘/dev/nrmt1h’ on Ultrix systems, and archive weekly files from sub-directory ‘91 50a’ and all requested reformatted data from both sub-directories.

IDL> mktap, 1, datadir=’91 50a,91 51a’, /alldata

same as above command, except that there will be no prompt for which reformatted files to place on the tape as all reformatted data from both directories will be archived.

As final comment there are a number of constraints one should be aware of when creating archive tapes. First, one must have WRITE privilege to the data directory. This is necessary since MKTAP creates a script file and two versions of the tape directory files in the data directory. Secondly, an observing log with the name which corresponds to the dates of the data archived must be on the directory or the program will prompt the user for its name.

B Institutes and Nodes with the Yohkoh Package

B.1 List of Institutes

The following are a list of the Yohkoh principle investigator and co-investigator sites which have the Yohkoh software and database package. The node names and a contact person are listed.

Institute

Computer Nodes Contact Person

Institute of Space and  isass0 thru isass4  Freeland
Astronautical Sciences (ISAS)  flare1 thru flare12  Shimizu

isasxa thru isasxc Bentley

isass6  Hudson

Lockheed Palo Alto Research sxt0 thru sxt6, sag Freeland Laboratory (LPARL)

Montana State University  sxt4, msusxt6  Acton
University of Hawaii  akala, akoko, maile  Canfield

hau, iliahi, wiliwili

University of California, Berkeley sunspot

McTiernan

Stanford University flare, panache Klimchuk Solar Physics Research noao

Harvey

Corporation (SPRC) National Astronautical Observatory spot

Tsuneta

of Japan (NAOJ) University of Tokyo neptune

Khan

Mullard Space Science MSSL Cluster Bentley Laborartory (MSSL) Rutherford Appleton solg1,RLSAC? Phillips Laboratory (RAL) Naval Research Laboratory (NRL) aspen,dogwood,sola Mariska

solb

This next table shows the various other institutions which are not primary or co-investigator institutions and which have the Yohkoh package.

University of Alabama at Huntsville nessie

Mariska (NRL)

University of Birmingham ??

Simnett

Solar Data Analysis Center (GSFC) umbra,sdac

Gurman,Zarro

California Institute of Technology sundog

Zirin,Lim

NSSDC (GSFC)

??

Batchelor

Marshall Space Flight Center golem, golem2 Porter University of Glasgow astro?

Norman?

Harvard-Smithsonian Center for darmok?

Kline?

Astrophysics National Center for Atmospheric riker

Stanger

Research (NCAR/HAO) Harvard University daystrom

Bookbinder

Stanford University  quake, kilohoku  Milford,Gethyn
Space Environment Laboratory, NOAA  seldmc,sunspot  McAllister,Bornmann

Department of Astronomy, Kyoto ??

Kawai

University Geophysical Institute, Kyoto step4ku

Yatagai

Unversity CRL - Akioka planet

Akioka

CRL - Marubashi?

??

??

Shibaura? Institute of Technology ogwl

Sabah

Nagoya ?  ??  Kato
Nobeyama ?  ??  ??

Astronomical Institute, Czech Republic 147.231.24.1

Farnik

B.2 List of Node Names and Numbers

A list of the most of the computer nodes which have the Yohkoh software is listed here. To verify that the lists are accurate, we recommend logging into sxt, isass0 or flare2 and looking

at the file /etc/hosts.
       DECNET             Internet                          Machine
SAG      24.131   192.68.162.134 (sag.space.lockheed.com)   DEC-VMS
SXT      24.134   192.68.162.107 (sxt.space.lockheed.com)   SGI-Unix
SXT1     24.362   192.68.162.100 (sxt1.space.lockheed.com)  DEC-Ultrix
SXT2     24.365   192.68.162.109 (sxt2.space.lockheed.com)  DEC-Ultrix
SXT3     24.366   192.68.162.110 (sxt3.space.lockheed.com)  DEC-Ultrix
SXT4              192.31.215.58  (sxt4.oscs.montana.edu)    DEC-Ultrix
SXT6              192.68.162.158 (sxt6.space.lockheed.com)  DEC-Ultrix
spot              133.40.2.120 (spot.mtk.nao.ac.jp)         SUN
neptune           133.11.3.23  (neptune.astron.s.u-tokyo.ac.jp) SUN

ISASS0   40.932   133.74.8.100 (isass0.solar.isas.ac.jp)    DEC-Ultrix
ISASS1   40.933   133.74.8.101 (isass1.solar.isas.ac.jp)    DEC-Ultrix
ISASS2   40.934   133.74.8.102 (isass2.solar.isas.ac.jp)    DEC-Ultrix
ISASS3   40.935   133.74.8.103 (isass3.solar.isas.ac.jp)    DEC-Ultrix
ISASS4   40.936   133.74.8.104 (isass4.solar.isas.ac.jp)    DEC-Ultrix
ISASS6            133.74.8.106 (isass6.solar.isas.ac.jp)    SUN
ISASXA   40.927   133.74.8.110 (isasxa.solar.isas.ac.jp)    DEC-VMS
ISASXB   40.928                                             DEC-VMS
ISASXC   40.929                                             DEC-VMS
FLARE1            133.74.8.1   (flare1.solar.isas.ac.jp)    SUN
FLARE2            133.74.8.7   (flare2.solar.isas.ac.jp)    SUN
FLARE3            133.74.8.3   (flare3.solar.isas.ac.jp)    SUN
FLARE4            133.74.8.4   (flare4.solar.isas.ac.jp)    SUN
FLARE5            133.74.12.94                              SUN
FLARE6            133.74.8.6   (flare6.solar.isas.ac.jp)    SUN
FLARE7            133.74.8.87  (flare7.solar.isas.ac.jp)    MIPS
FLARE8            133.74.8.88  (flare8.solar.isas.ac.jp)    MIPS
FLARE9            133.74.8.89  (flare9.solar.isas.ac.jp)    MIPS
FLARE10           133.74.8.90  (flare10.solar.isas.ac.jp)   MIPS
FLARE11           133.74.8.91  (flare11.solar.isas.ac.jp)   MIPS
FLARE12           133.74.8.92  (flare12.solar.isas.ac.jp)   MIPS
FLARE13           133.74.8.93  (flare13.solar.isas.ac.jp)   MIPS
FLARE14           133.74.8.94  (flare14.solar.isas.ac.jp)   MIPS
FLARE15           133.74.8.95  (flare15.solar.isas.ac.jp)   MIPS
MSSL     19.253                                             DEC-VMS
MSSLV1            128.40.70.62 (msslv1.mssl.ucl.ac.uk)      VMS-4000
MSSLV2            128.40.70.63 (msslv2.mssl.ucl.ac.uk)      VMS-4000
MSSLVB            128.40.70.70 (msslvb.mssl.ucl.ac.uk)      VMS-3100
MSSLA1            128.40.70.58 (mssla1.mssl.ucl.ac.uk)      VMS-3000
ssd0              128.60.8.1    (ssd0.nrl.navy.mil)         VAX
aspen             128.60.8.72   (aspen.nrl.navy.mil)        SUN
dogwood           128.60.108.102 (dogwood.nrl.navy.mil)     SUN
sola              128.60.8.41   (sola.nrl.navy.mil)         SUN
solb              128.60.8.42   (solb.nrl.navy.mil)         SUN
nessie            146.229.9.11  (nessie.optics.uah.edu)     SUN
koa               128.171.167.1 (koa.ifa.hawaii.edu)        SUN?
mamane            128.171.2.2   (mamane.ifa.hawaii.edu)     SUN

sunspot                         (sunspot.ssl.berkeley.edu)  SUN?
star              36.10.0.5     (star.stanford.edu)         ???
flare             36.14.0.166   (flare.stanford.edu)        DEC-Ultrix
panache           36.14.0.51    (panache.stanford.edu)      DEC-Ultrix
ogwl                            (ogwl.se.shibaura-it.ac.jp) ???
setp4ku                         (step4ku.kugi.kyoto-u.ac.jp) SUN
riker                           (riker.hao.ucar.edu)        SGI
darmok                          (darmok.harvard.edu)        ???
astro                           (astro.glasgow.ac.uk)       ???
sundog   14.905   131.215.139.154 (sundog.caltech.edu)      VMS
umbra             128.183.57.24 (UMBRA.gsfc.nasa.gov)       DEC-Ultrix
NOTE: To go from DECNET xx.yyy convention to the single node number
notation, take the xx portion, multiply by 1024, and add `yyy'.
So SAG becomes 24*1024 + 131 = 24707::

C Accessing the University of Hawaii SPAM Database

Mees Solar Observatory (MSO) is located on Haleakala, Maui, at 165E longitude and 20N latitude. MSO instrument operation hours vary from 1800-0300 UT in December to 1630-0400 UT in June. A searchable catalog called SPAM (Spectroscopy and Polarimetry at Mees) contains MSO data logs (not data). Some MSO data are available on line. Other MSO data can be obtained through the Data Use Coordinator, canfield@mamane.ifa.hawaii.edu.

SPAM logs (and searchable attributes, in addition to date, day of year, and time) include: MSO Log (instrument, NOAA AR, data type, observing setup), SEL Event List (NOAA AR, X-ray Class), SEL Region Report (NOAA AR), Yohkoh Orbit Summary, Yohkoh SXT Quiet Mode PFI Observations (latitude, longitude, X-ray and optical image size), Yohkoh Flare Observations (latitude, longitude, specific channel counts or ratios). SPAM runs on a Sun Unix workstation at MSO, and is available over Internet. Simply access (e.g., telnet) koa.ifa.hawaii.edu (128.171.167.1) from any vt100, Sun, or xterm emulator. Log on as spam (lower case); there is no password. New users are asked to read release notes and hints.

Vector magnetograms from the MSO Stokes Polarimeter are available on line for about a month after acquisition. These data are available by anonymous ftp over Internet. Use ftp koa.ifa.hawaii.edu, log in as anonymous, provide your name as the password, and then respond to the ftp prompt with cd pub/stokes. Use get to obtain a copy of the README file for further information and conditions for data use.

D Accessing the Optical Solar Data of NAOJ

1. Introduction
This document describes how to make access to various optical data of
the sun obtained at the National Astronomical Observatory of Japan (NAOJ).
>From workstations which support the X-window and are connected to NAOJ via
network, you may log-in to our workstation
     'spot' (spot.mtk.nao.ac.jp, IP-address=133.40.2.120),
under the username 'guest'. The password can be obtained by sending
an e-mail request to either one of the following.
     sakurai@spot.mtk.nao.ac.jp (T.Sakurai)
     ichimoto@spot.mtk.nao.ac.jp (K.Ichimoto)
For remote users it would be more convenient to transfer the data and
the software to your own workstation and run the software locally.
The IDL package should be installed on your workstation. Files can
be transferred by anonymous ftp on 'spot'. After you are connected to
'spot' via ftp, enter 'anonymous' as the username. For password prompt,
you are requested to enter your name, which will be recorded.
These data can be utilized freely in your data analysis. If you write
a paper or give a talk by using these data, you are requested to consult
with us in terms of co-authorship.
For technical details, please contact:
Norikura heliograms/filtergrams          : K.Ichimoto, K.Kumagai
Flare Telescope magnetograms/Dopplergrams: K.Ichimoto, T.Sakurai
Okayama magnetograms                     : T.Sakurai
H-alpha full-disk images                 : Y.Suematsu, N.Tanaka
These people can be accessed by e-mail at <name>@spot.mtk.nao.ac.jp.
In any case, only representative data (a few images per day)
are provided on-line. Once you figure out that the data you are
interested in are in archive tapes, you are encouraged to visit
NAOJ, recover the data from tapes, and make use of them.

2. Observing Instruments
This section briefly describes the sources of data, namely the optical
observing facilities currently being operated at NAOJ.
2.1  Mitaka Campus
(a) H-alpha flare patrol
optics: 4cm refractor + Halle H-alpha filter
field of view: full disk
pixel size: 4"
form of data: digital images of H-alpha line center
image cadence: one minute
recording media: digital audio tape
(b) The Solar Flare Telescope
This is made of four telescopes T1-T4.
field of view: 340" x 320"
pixel size: 0.66"
(T1) vector magnetograph
optics: 20cm refractor + birefringent filter (band width=1/8A)
form of data: digital I,Q,U,V intensities of Fe I 6303A line
image cadence: 3 minutes
recording media: digital audio tape
(T2) continuum telescope
optics: 15cm refractor + broad-band filters
form of data: analog images taken mostly at the wavelength of g-band
image cadence: video rate
recording media: SVHS video tape
(tapes are recycled unless no notable events are found)
(T3) H-alpha telescope
optics: 15cm refractor + Zeiss H-alpha filter
form of data: analog images of H-alpha line center
image cadence: 10 seconds
recording media: laser video disk
(T4) Dopplergraph
optics: 20cm refractor + birefringent filter (band width=1/5A)
form of data: digital I,V,Doppler shifts of Fe I 6337A line
image cadence: 3 minutes
recording media: digital audio tape
(c) STEP Full-Disk Magnetograph (STEP=Solar Terrestrial Energy Program)

20cm heliostat + 6cm refractor + birefringent filter
field of view: full disk
pixel size: 5"
form of data: digital I,V,Doppler shifts of Fe I 5324A line
image cadence: (the instrument is now under testing)
recording media: digital audio tape
2.2  Norikura Coronagraph Station
(a) Automated Digital Coronagraph
optics: 10cm coronagraph + interference filters
field of view: 2600" x 2400"
pixel size: 5"
form of data: digital images of the corona in the following wave bands
     wavelength               band width
     5303A                    3A
     He D3                    4A
     H-alpha                  3A
     continuum at 6630A       21A
image cadence: 3 minutes
recording media: digital audio tape
(b) Spectroheliographic Observations
optics: 25cm coronagraph + Littrow spectrograph
raster area: four swaths of 500" wide (2.5" step)
             synthesis of full disk maps is made after the observation
spectral coverage: 2A (0.2A x 10 wavelength points)
form of data: digital spectroheliogram of He I 10830A
image cadence: one map per day (one raster scan takes 30 minutes)
recording media: optical disk
2.3 Okayama Observatory
optics: 65cm reflector + Littrow spectrograph + photoelectric magnetograph
raster area: 500" x 450" (10" step) or 420" x 390" (6" step)
spectral coverage: 27-80mA from the center of 5250A line
           (detectors are photomultipliers; no wavelength resolution)
form of data: digital I,Q,U,V,Doppler shifts of Fe I 5250A
image cadence: a few maps per day (one raster scan takes one hour)
recording media: magnetic tape and magneto-optical disk
The data available on-line are:
part of Mitaka H-alpha full-disk images (2.1a),

part of Flare Telescope vector magnetograms (2.1b-T1),
all of Norikura He 10830 fill-disk heliograms (2.2b), and
all of Okayama vector magnetograms (2.3).
Other data are archived in tapes in digital form, or recorded on analog media.

3. Browsing the Data by 'show'
After you log-in at 'spot' under the username 'guest', follow the instruction
shown on the banner message.  You simply type
spot% show
and then enter your machine's IP-address after the prompt. An interactive,
menu-driven session on the X-window will start.
By using the mouse you first select from
      magnetic data of flare telescope
      Ha full disk image of Mitaka
      Norikura data (He10830, corona)
      Okayama magnetic data
      exit
On all the menus, <return> means that you go back to the main menu, and
the 'show' program stops by selecting <exit>.
3.1 Flare Telescope Vector Magnetograms
The data have names 'fbyymmdd.hhmm', where yymmdd and hhmm designate,
respectively, the date and time in UT. By using the mouse, select the file
you want to see. When the menu is too big and is split into several pages,
you may select 'next page' and the next page of the file list will
be shown. The other options control plotting styles and printer output.
<contour> : magnetogram in contour
<spot+contour> : magnetogram in contour, sunspots in gray scale
<red-blue> : magnetogram in red-blue, sunspots in contour
<PostScript> : make postscript output file
<printout> : print postscript file at a local printer at Mitaka
<PostScript> is only selectable after you have selected data to be displayed.
In order to transfer the postscipt file, use anonymous ftp. The contour plot
is saved in 'tmp/magcont.ps' under the anonymous ftp directory.
After you exit, the data are stored in arrays: intensity map in 'img',
longitudinal magnetid field in 'Bl', and transverse magnetic fields in

'Bx,By'. The header h contains observing parameters. The magnetic data
are 2 byte integer in unit of 0.1 Gauss and each image has 256*240 pixels,
where a pixel corresponds to about 1.35 arcseconds.  The instrument tends to
saturate for magnetic fields stronger than about 1700 Gauss. The field
strength in large sunspots should be used with some caution.
The direction of the transverse magnetic field (i.e. the resolution of
the 180 degree ambiguity) is determined by referring to the potential field
calculated from the longitudinal magnetic field distribution. The direction
may not be correct in regions with high magnetic shear.
In order to obtain the observation status, type
IDL> fstplot
and then choose yymmdd.
3.2 Mitaka Full-Disk H-alpha Images
The data have names 'hayymmdd.hhmm', where yymmdd and hhmm designate,
respectively, the date and time in UT. By using the mouse, select the file
you want to see. After you exit, the image data are stored in an array
'img', and the header h includes the observing parameters.
3.3 Norikura Spectroheliograms
In the 'Norikura' menu ( He10830, corona ) you should select He10830.
The other entry is not yet implemented.  The data have names 'heyymmdd.hhmm',
where yymmdd and hhmm designate, respectively, the date and time in UT.
The data names followed by '.v' are Doppler images. By using the mouse, select
the file you want to see. After you exit, the image data are stored in an array
'img', and the header h includes the observing parameters.
3.4 Okayama Vector Magnetograms
First you select the year (1982 - ), and the file list appears.
The data have names 'ymddn.D', where y is a letter indicating the year
(B=1982, C=1983, ,,, M=1993), m indicates the month (1=Jan, ,, 9=Sept,
A=Oct, B=Nov, C=Dec), dd is the date, and n=A,B,C,,, are sequence
numbers of observation on the day. Then you are asked to enter
several numbers from the keyboard as follows.
obs.area    : 0=skip, 1=show: _
     (location of the observed area on the sun)

halftone map: 0=skip, 1=    I, 2=   BL, 3=   VL: _
contour map : 0=skip, 1=    I, 2=   BL, 3=   VL: _
vector map  : 0=skip, 1=(   BX,   BY): _
Here I is the intensity map, BL is the longitudinal magnetogram, VL
is the Doppler map, and (BX,BY) are the transverse field vectors. If 0 is
entered, the corresponding map is not shown. If the potential field lines
had been computed and stored in a file 'ymddn.L', you will be asked:
field lines : 0=skip, 1=show : _
At the next prompt
1=plot again, 2=hardcopy, 9=exit > : _
you can plot the map in a different style by selecting 1. The postscript
output file will be created if you select 2. The postscript file can
be later transferred by anonymous ftp.
After you exit, the data are stored in a floating array 'rdata',
and the structure 'mparam' contains observing parameters.
The 'mparam.status' includes fields such as
i, q,u,v, bx,by,bl,,,,.
If for example the data By are stored in rdata(*,*,2), mparam.status.by
takes the value 2. In order to know which part of rdata contain valid
data, type
IDL> help,/st,mparam.status
The data whose status value is -1 are not stored in rdata array.
The electric current distribution can be seen by typing
IDL> .run jn
IDL> magplot, mparam, rdata, /nomenu
Then you will see the following prompt.
obs.area    : 0=skip, 1=show: _
halftone map: 0=skip, 1=    I, 2=  BLP, 3=   BL, 4=   VL, 5=   JN: _
contour map : 0=skip, 1=    I, 2=  BLP, 3=   BL, 4=   VL, 5=   JN: _
vector map  : 0=skip, 1=(  BXP,  BYP), 2=(   BX,   BY): _
1=plot again, 2=hardcopy, 9=exit > : _
Here JN is the normal component of the electric currents, and the field
components followed by P are those of the potential magnetic field.

4. Anonymous ftp Access to spot

When you are connected to 'spot' via anonymous ftp, under the home directory
you have directories
flare                           : Flare Telescope magnetograms
         - log                  : Flare Telescope magnetograms
monochro                        : H-alpha full-disk images
norikura - he10830              : He10830 spectroheliograms
                 - corona
okayama  - 1982                 : Okayama magnetograms
                 - ...
doc                             : documents
idlsoft                         : IDL routines
All the IDL routines used here are stored under 'idlsoft'.
Compressed tar files of IDL software is provided as naojidl.tar.Z
in the home directory. The 'doc' directory contains text files explaining
the software and the data analysis.
The observation log files of the Flare Telescope are in /solar/flare/log.
The files 'yymmdd.log' give you the time and the observing region
for all data on the day. The files 'lv##.log' contain list of H-alpha
video disk recordings of the Flare Telescope T3.

5. Description of the Software
The 'show' program assumes the directory structure of NAOJ workstations.
Therefore, when you want to display and analyse the data obtained via ftp,
you have to invoke lower-level IDL routines.
5.1 Flare Telescope Vector Magnetograms
To read a file named 'fbyymmdd.hhmm' and store the data into img,Bx,By,Bl:
IDL> f1data, 'fbyymmdd.hhmm', h, img, Bx, By, Bl
Or an interactive program can be invoked by
IDL> .run fdisp
The map can be displayed by
IDL> magtv1, h=h, img, Bx, By, Bl
or
IDL> magtv2, h=h, img, Bx, By, Bl

or
IDL> magcont, h=h, img, Bx, By, Bl
To calculate and display electric current distribution:
IDL> .run jnft1
(The data should exist in h,img,Bx,By,Bl.)
5.2 Mitaka H-alpha Images and Norikura He10830 Heliograms
To read a file and store the image into 'img':
IDL> nkrget, <filename>, h, img
To display:
IDL> tvscl, img
5.3 Okayama Vector Magnetograms
To read a magnetogram file:
IDL> readm, mparam, rdata, filename=<filename>
To display magnetograms and potential field lines:
IDL> magplot, mparam, rdata, /nomenu
To calculate and display electric current distribution:
(The data should exist in mparam and rdata.)
IDL> .run jn

E Using the Laser Disk System at ISAS

E.1 Viewing Existing Movies (VIDEO MENU)

This program provides interactive computer controlled viewing of the library of movie sequences recorded on the ISAS SONY laser disks. The user must be logged on to the isass3 computer (remote logon from another ISAS workstation is fine) in order to use the SONY laser disk player.

To run the software the user must log on to isass3. The calling sequence is:

IDL> video menu

The user is prompted to check for the proper settings on the SONY disk drive, and then a listing is displayed from which the user selects a movie sequence for display. Upon selection the sequence is played repeatedly. While the sequence is playing the user may adjust the speed by keyboard input. The user may interrupt the movie at any time and return to the menu. The library of available movie sequences is updated manually as movies are recorded on the SONY, so the listing could be out of date.

In addition to this basic mode, VIDEO MENU may be called with a keyword vcr set which allows non-interactive playing of frame sequences specified in an input file. This mode of operation is convenient for demonstrations and for making video tape copies of selected sequences of movies.

The centerpiece of the laser disk movie catalogue is the full disk movie which covers the duration of the mission. However, there are over 100 entries in the menu, corresponding to a variety of interesting sequences showing flares, large scale coronal transients, and theoretical modeling.

E.2 Making Title Pages

The video title pages may be prepared by using an IDL program called VIDEO TITLE. First, you must prepare an ASCII control file that specifies what the title pages should contain. The IDL routine reads this file and writes the text to a Z-buffer window (using ordinary XYOUTS commands) and then reads it back into an array that can be stored in a *.genx file. The font type and size and placement can be easily adjusted. Some example calls are:

IDL> array = video title(filename) IDL> array = video title(filename,/plot) ; Display the results IDL> array = video title(filename,/plot,outfile=’video1’,/high) IDL> array = video title(filename,/plot,outfile=’video1’)

The variable filename specifies the name of an ASCII control file that contains the text of the page. For example try the following:

IDL> array = video title(’/ys/gen/doc/howto video title.doc’,/plot)

where /ys/gen/doc/howto video title.doc is the name of a “how to” file which is also an example control file. This file explains the format of the control file. For the first time users, it is recommended to copy that file over and make changes to it to accommodate your needs.

The control file specifies the position on the screen, the text, the size and thickness of the characters. Different fonts can be used also. A portion of that file is shown here:

;  howto_video_title.doc                J. R. Lemen  Rev A,  4-Jan-94
;
;  The control file contains 5 items per line separated by a double
;  slash or "//".   The first and second fields are the positions in
;  the 620 by 480 sized video window (0,0 is the lower left corner).
;  It is advisable not to write too close to any edge, since the
;  text may be lost, depending on the display monitor.  Setting xpos
;  equal to -1 will cause the text to be automatically centered (and
;  leading or trailing blanks will be discarded).  The third field
;  is the text and the fourth and fifth fields are the character
;  size and thickness, respectively.
;
;  The Fonts can be changed using the standard !n IDL vector commands
;  (see the example below).
;
;  This entire document is an example control file (since every line
;  except the important ones below begin with an ";").  Below the
;  dashed line is a specific example.  This file will cause two
;  pages to be generated (because of the NEWPAGE// line).  It is not
;  necessary to have more than one title page per file.
;
; ------------------------------------------------------------------
;  Semicolons are treated as comments
;  Separate fields with //
;  xpos // ypos // Text // size // thickness
;  (size and thickness will default to 1.0 if omitted)
;  (setting xpos to -1 will cause x position to auto-center)
-1//425//X-RAY CORONA OF THE SUN                           //2.8//2.8
-1//375//Images from the Soft X-ray Telescope              //2.1//1.4
-1//345//on the YOHKOH mission                             //2.1//1.4
-1//275//Data from:  November 11 to 27, 1991               //2.1//1.4
-1//220//!8Prepared by:                                    //2.1//1.4
-1//190//J. Lemen, G. Slater, K. Strong, and S. Tsuneta    //1.7//1.7
-1//100//Institute for Astronautical Research, Japan       //1.4/1.4
-1// 75//National Aeronautics and Space Administration, USA//1.4//1.4

-1// 30//(V1.0 13 April 1992)//1.4
NEWPAGE//
-1//300//Data from:  November 11 to 27, 1991               //2.1//1.4
-1//200//!18Displayed 100,000 times real time              //2.1//1.4

The /high switch makes the routine run slow, but gives a higher quality result. It is recommended to debug the output without the /high switch and only use it for the final result.

The outfile = switch will cause the title pages to be written to a *.genx file that can be written to the ISAS Sony Laser disk using GO NVS5. If you want the background color to be blue, make sure that you specify color table 1 in GO NVS5.

E.3 Recording a New Movie to Laser Disk

E.3.1 Hardware: Positioning the Sony Laser Disk

1. Check that the green light on the NVS is on. Power cycle if necessary.

2. Check that the RS-232 cable is connected from the NVS to the Sony LVR-5000.

3. Make sure the correct optical disk is inserted (check side A or B).

4. Put the ‘LOCAL/REMOTE’ switch in the local position.

5. Depress the ‘REC STANDBY’ button to initiate blank frame check. This will find positions of all blank areas on the disk. When completed, the disk will be positioned to the last recorded frame before the nearest (not necessary the first) blank frame.

6. Use the record ‘SKIP’ buttons to position the laser disk to the last recorded frame before the desired starting frame number (consult the Sony Laser disk log book). On the display, Cont. Rec = is the last recorded frame before the blank frame; DUR = number of blank frames.

7. Toggle the format of the numbers displayed to show the time-format. Record numbers in the log book. To start over, press the ‘CLEAR’ button.

8. Set the left console switch to ‘remote’.

E.3.2 Software

1. Login onto a non-isass0 machine (isass1-4).

2. Enter IDL. Make sure no other IDL window jobs are running on the same workstation. If there are two jobs, the color table could be corrupted.

3. Then enter one of the following:

IDL> go nvs5,/sfd ; Standard weekly movie making IDL> go nvs5,/sfd,/nosfw ; Standard weekly movie making IDL> go nvs5,/help ; For on-line documentation IDL> go nvs5,/genx ; to use a GENX type file as the input

The temporary NVS files will be written to your home directory. Use the nvsdirfn=’/1p/user/nvs0’ switch to redirect to another location.

For the SFD option, the default increment is 10 (process 10 images at a time). This is good if no one else is using the workstation at the same time. If the /sfd switch is given, go nvs5 will prompt the user with the list of files which are in the /yd1/sfd directory.

For the GENX option, the array will not be autoscaled (it should be a byte type). If the number of images in the genx file is large, IDL may run out of core. In that case, try increment=1. The program will prompt the user for the GENX filename.

E.3.3 Adding Another Movie

The above procedure is for starting from scratch. If you have just completed a run and the Sony frame number has not been changed, then just type:

IDL> go nvs5 [,/sfd]

to make the next movie.

E.3.4 NVS/Sony Problems

• Invalid Num – If the message ‘Invalid Num’ appears on the display monitor, the nvs box is not in sync with the Sony laser disk. To correct this situation begin with a blank frame check (see above hardware procedure) and follow that with software procedure shown above.

• One common problem is to forget to attach the RS-232 cable. This is sometimes switched to ISASS3 to control the recorder during playback to record with the VCR).

• How to toggle the format frame/time format: depress the frame/time button located on the lower-left side of the console number keypad–where frame # is a simple number and time has the format of mm:ss:ff.

E.4 Transferring a Movie to VCR Video

MK VCR is an IDL procedure which can be used to control the Sony Laserdisk recorder. In this way sequences can be displayed easily in order to make VCR video recordings.

The MK VCR procedure reads an ASCII file (described) below and then calls sonyloop to actually control the laser disk player.

E.4.1 Initial setup

1. First, make sure the RS-232 cable from isass3 is connected to the back of the Sony recorder (it requires an adapter connector).

2. Switch the local/remote switch to remote.

3. Log on to isass3 and enter IDL.

E.4.2 Short test of hardware communications.

You can preform a simple test to see if things are working correctly by typing:

IDL> sonyloop,6000 ; Position laser disk to frame number 6000 IDL> sonyloop,6000,6200 ; Play from 6000 to 6200 at normal speed IDL> sonyloop,6000,6200,sec=2 ; Play a factor of 2 slower

E.4.3 Running MK VCR

If things look like they are working correctly, you can then try mk vcr. To get a template of the file the mk vcr will read, execute the following Unix command to copy the example file to your own directory:

% cp /ys/gen/data/mk vcr example.txt .

After you copy the example file, do the following. Insert Side B of the laser optical disk. Then,

IDL> mk vcr,’mk vcr example.txt’,1,3

which execute lines 1 to 3 of the file. (Line 0 puts up a blank frame). If you type

IDL> mk vcr,’mk vcr example.txt

the routine will inform you of the number of lines in the file and ask you to enter the start and stop line numbers.

If you want to make a movie, you could try the following IDL program:

IDL> sonyloop,8000 ;Get it to a blank frame IDL> for i=0,2 do mk vcr,’mk vcr example.txt’,0,23 ;There are 24 lines in the file IDL> sonyloop,8000 ;End on a blank frame IDL> end

As you can see, the sequence in the ‘for loop’ will execute 3 times. Each time through takes about 7 minutes, so the whole movie will be just over 20 minutes in length.

E.4.4 How to create/modify the file that MK VCR uses

The file has some descriptions about how to change it. Basically, the lines in the file look like the following:

 8000 ; Position to a blank frame (provide leader)
 6077, 6077,    4,    0, [ -1  0  0  0] "Movie Title Page"
 6078, 6078,    2,    1, [ -1  0  0  0] "Nov 1991 Movie - Title"
 5048, 5225,    2,    1, [  1  2  1 -1] "Nov 1991 Movie"
; c0    c1      c2    c3       [c4]         "  c5   "

The above sequence will display the Nov 91 press release sequence if Side B is inserted. Some notes about the format:

• Text following a ‘:’ is considered to be a comment.

• Only the first column is required. All subsequent columns are optional.

• c0 is starting frame number.

• c1 is the ending frame number.

• c2 is the wait (in sec) after displaying the first frame

• c3 is the wait (in sec) after displaying the last frame

• c4: The number in the boxes are the speeds that are used by sonyloop. Up to four different speeds can be specified.

      0 is no operation
      1 is normal speed (30 frames per second)
      2 is twice slower (15 frames per second)
      3 is 3x slower, etc (10 frames per second)
      -2 is 3x faster.

If a fifth parameter is specified in the brackets, it is a repeat number. If this number is negative, sonyloop will play the sequence forward and backward. For example,

 5048, 5225, 2,  1, [  1  0  0  0  1] "Nov 1991 Movie" ;1 time forward
 5048, 5225, 2,  1, [  1  0  0  0 -1] "Nov 1991 Movie" ;1 time forward/backward
 5048, 5225, 2,  1, [  1  0  0  0 -3] "Nov 1991 Movie" ;3 times forward/backward

SPECIAL NOTES:

• Sonyloop doesn’t send the correct commands to the Sony Laser Recorder to go backwards. Thus, no matter what speed you specify, it always goes at normal speed in reverse. (Maybe a Japanese reader can study the manual and fix this problem).

• Sonyloop tries to compute the time it will take the laser disk recorder to play the specified sequence. Because the timings in IDL can be off by a certain amount (less than or equal to about a second), it is wise to make one of the wait parameters (c2 or c3) equal to at least 1. If they are set to 0, IDL will sometimes send the next sequence before the previous sequence has had a chance to complete.

F Using the Laser Disk System at LPARL

The following is a guide to record and play laser disk movies at LPARL.

F.1 Viewing Existing Movies

F.1.1 Using IDL and Menu Driven Software

The program VIDEO MENU described in section E.1 is not available at LPARL during the writing of this guide, however it may exist now. See one of the software engineers to see if it is available.

F.1.2 Displaying a Single Recorded Image

A. Manual Operation

The Panasonic optical disk recorder/player and the Panasonic optical disk player may be “manually” controlled by using either the buttons on the front of each machine or the remote controllers. These buttons and the remote controller are labeled in English and are easily understood.

B. IDL Operation

1. Log onto sxt0 on any terminal

2. Enter IDL

3. Suppose you want to display frame 1000. Here’s the command:

IDL> pana,1000

F.1.3 Playing the Panasonic Laser Disks

A. Manual Operation

The Panasonic optical disk recorder/player and the Panasonic optical disk player may be played “manually” by using either the buttons on the front of each machine or the remote controllers. These buttons and the remote controller are labeled in English and are easily understood.

B. IDL Operation

1. Log onto sxt0 on any terminal

2. Enter IDL 3. Suppose you want to play frames 100 to 200 on the recorder for 10 repetitions at a rate

of 90 frames per second. Here’s the command:

IDL> movie pana,100,200,rate=10,num repeat=3

4. The general command is

IDL> movie pana,start frame,stop frame,device,device=device1,rate=rate, $ IDL> num repeat=num repeat,backforth=backforth

The device parameters should be ignored because at the present time there is only one available device, which is the default. More information on this routine can be obtained with the command:

IDL> doc library,’movie pana’

F.2 Making Title Pages

See the description in the previous chapter on page 90 describing the laser disk system at ISAS.

F.3 Recording a New Movie to Laser Disk

1. Turn on power to Panasonic recorder and Sony monitor labeled ‘Laser Disk Recorder’.

2. Find the disk you want to record on (it may be in the recorder). Note that each disk has two sides. The ‘up’ side is the active side. Note also that there is blue ‘write protect’ sliding switch on the right side of each side of the disk. Move this switch to its right hand position (‘read/write’) using a Bic pen or equivalent tool. Now insert the laser disk you want to record on into the recorder after first ejecting the disk already in there if there is one.

3. Log onto sxt0 on an X terminal

4. Enter IDL

5. At the prompt do this:

IDL> ptest

Look at the monitor. If after a few seconds a test pattern appears the equipment is properly set up and you may proceed. If no test pattern appears or if it looks unusual then something is wrong. Refer to section IV on problems below.

6. Images may be recorded either from an SDA file or from a data cube resident in memory. The images will be re-sized to fit within the 630 x 486 size of the Peritek video image. The re-sizing will use linear interpolation unless the /NOINTERP option is used, By default they will be centered. The images may be of any size. If they are in a data cube then the entire cube will be normalized to its maximum value and byte scaled, unless the /NOSCALE keyword is used in the procedure call. If the images are in an SDA file the images must be pre-scaled by the user. Decide on a color table. A gray scale (IDL standard color table 1) will be used by default. If the color table desired is not one of the first 15 IDL standard color tables then you must supply r, g, and b color vectors.

7. Decide what frame on the laser disk you wish to start recording at. The IDL recording program will automatically find the first sufficiently large unrecorded area if you skip this step. To do this manually, type:

IDL> command pana,’OF’

Type the ‘display’ button on the remote controller to display the current frame number on the monitor. Then use the remote controller to play or scan through the disk until you find the frame number at which you want to start recording. Key the number of frames you wish to record using the keypad. Then hit the red ‘rec mode’ key. This will verify that there are enough contiguous blank frames available following the selected frame.

8. Suppose you have the cube in a 3-D IDL array called DATA and the color vectors in the variables R,G, and B. Here’s how to record it:

IDL> pana record,data,r=r,g=g,b=b,/record

9. For recording images from an SDA file the call would be

IDL> pana record,file name,r=r,g=g,b=b,/record

where file name is a string naming the SDA file, including path. More information on this routine can be obtained like this:

IDL> doc library,’pana record’ F.3.1 Panasonic Problems

If no test image appears after issuing the PTEST command in the instructions described in the previous section, then do these things:

• Make sure that the power switch on the back of the Peritek ‘clock box’ is in the on position. The clock box is located in the middle shelf of the SXT0 computer cabinet.

• Check that the four gold colored video cables from the Peritek are connected to the RGB and sync terminals on the back of the recorder.

F.4 Transferring a Movie to VCR Video

As of the writing of this document, the routine MK VCR (which is described on page 93) was not available for the LPARL video system. The routine may be expanded in the future. The best method for transferring a laser disk movie to VCR video is to talk to one of the LPARL software engineers.

G Using Magneto-Optical (MO) Disks at ISAS

A complete archive of the Yohkoh data for the whole mission is available on Magneto-Optical (MO) disks at ISAS.

A Unix script and associated alias is available to mount and dismount MO (magneto optical) disks at ISAS. An IDL interface routine is also provided to allow MO mounts and dismounts from within an IDL session or callable from other IDL routines. Once mounted, access to data on these drives is identical to other magnetic disks — standard Unix operations (ex: ls) are permitted. For MO disks containing Yohkoh data, the standard IDL functions (ex: yodat access) are available.

MO disks are available on several machines of the SUN machines at ISAS. including flare3, flare4, flare6, flare13, flare14, flare15. The first (and in most case, only) MO drive on any machine is referred to as drive 0. Some machines may have a second drive (drive 1) — all drives are clearly marked with a drive number. Find an available (empty) drive and insert your disk — if you intend to write, verify that the write protect switch allows writing. By default, the mount is executed on the machine which contains the MO driver. If that machine is different than the machine which executes the command, then the MO disk is automatically ’exported’ and mounted on the local machine. Use the /all switch to force mounting on all ISAS machines.

G.1 Mounting MO Disks From Unix

The alias momount executes the script which performs all MO operations. Parameters and switches allow specification of machine name, drive number, access mode, dismounts, etc. Options include /write /umount /eject /1 and /all. Some sample calling sequences are:

% momount flare13 # mount flare13 , drive 0, readonly % momount flare13 /umount # dismount same % momount flare4 /1 /write # mount flare4, drive 1, read/write % momount flare4 /1 /eject # dismount and eject same % momount flare15 /all # mount flare15, drive 0, on all nodes

G.2 Mounting MO Disks From IDL

An IDL driver sets up the parameters and spawns a call to the same script file described above to allow IDL MO control (interactive or via routine). Keyword options similar to the Unix script are available with the IDL routine. Some examples are:

IDL> momount,’flare13’ ; mount flare13 , drive 0, readonly IDL> momount,’flare13’,/umount ; dismount same IDL> momount,’flare4’, /one, /write ; mount flare4, drive 1, read/write IDL> momount,’flare4’, /one, /eject ; dismount and eject same

H Guides to Yohkoh Operations at ISAS

The following documents are useful for understanding different aspects of the operations of Yohkoh.

/ys/gen/doc/ssoc tohban manual * A detailed description of how to perform

the SSOC tohban duties

auto toban2.pro The header of this program describes in detail the different steps on how the daily KSC reformatting is performed

/ys/gen/doc/ref super.doc This document describes the different steps which are performed during the standard reformatting

/ys/gen/doc/cron.doc Describes the different cron jobs that are run automatically at ISAS and LPARL to process data, distribute software and data, and to copy data from GBO sites

/ys/gen/doc/howto term images.doc Explains the procedure SSOC Tohbans should follow when planning SXT terminator images

/ys/gen/doc/krasm.doc Explains how the SSOC Tohbans may access the RASM scheduling computer by modem